### v1

# Simplicial cell decompositions of $\mathbb{CP}^{\hskip 0.85358ptn}$

Abstract
According to a well-known result in geometric topology, we have $\left(\mathbb{S}^{2}\right)^{n}\!\!/\operatorname{Sym}(n)=\mathbb{CP}^{n}$, where $\operatorname{Sym}(n)$ acts on $\left(\mathbb{S}^{2}\right)^{n}$ by coordinate permutation. We use this fact to explicitly construct a regular simplicial cell decomposition of $\mathbb{CP}^{n}$ for each $n\geq 2$. In more detail, we take the standard two triangle crystallisation $S^{2}_{3}$ of the $2$-sphere $\mathbb{S}^{2}$, in its $n$-fold Cartesian product. We then simplicially subdivide, and prove that naively taking the $\operatorname{Sym}(n)$ quotient yields a simplicial cell decomposition of $\mathbb{CP}^{n}$. Taking the first derived subdivision of this cell complex produces a triangulation of $\mathbb{CP}^{n}$. To the best of our knowledge, this is the first explicit description of triangulations of $\mathbb{CP}^{n}$ for $n\geq 4$.
MSC 2020: 57Q15; 57Q05; 05C15; 06A06.
Keywords: Complex projective spaces, triangulations of manifolds, simplicial cell complexes.
1 Introduction
A long-standing problem in combinatorial and geometric topology is to find an explicit triangulation of the $n$-dimensional complex projective space $\mathbb{CP}^{n}$ using as few vertices as possible. In [AM1991], Arnoux and Marin proved that any triangulation of $\mathbb{CP}^{n}$ requires at least $(n+1)^{2}$ vertices, and equality can only hold for $n\leq 2$. Of course, $\mathbb{CP}^{1}\cong\mathbb{S}^{2}$ has a unique four-vertex triangulation $S^{2}_{4}$, the boundary complex of the three dimensional simplex. Also, it has long been known that $\mathbb{CP}^{2}$ has a unique 9-vertex triangulation $\mathbb{CP}^{2}_{9}$, first found by Kühnel and Banchoff (see [AM1991, BD1994, KB1983]). But the problem of constructing triangulations as close to the Arnoux-Marin bound as possible remains open. Even more, for $n\geq 4$, no explicit triangulation of $\mathbb{CP}^{n}$ appears to be known. A representation of $\mathbb{CP}^{n}$ as a quotient of a torus by $\operatorname{Sym}(n+1)$ was used in [AM1991, § 14] to obtain an $(n+1)^{2}$ vertex weak cellulation of $\mathbb{CP}^{n}$. The authors of [AM1991] observed that, when $n\geq 3$, this weak cellulation can not be triangulated without adding extra vertices.
One idea to obtain reasonably small triangulations of $\mathbb{CP}^{n}$ is based on the following observation: There exists a natural (non-free) action of the symmetric group $\operatorname{Sym}(n)$ on $(\mathbb{S}^{2})^{n}$, the $n$-th Cartesian power of the $2$-sphere, by coordinate permutation. It is well-known that the quotient space $(\mathbb{S}^{2})^{n}/\operatorname{Sym}(n)$ is the $n$-dimensional complex projective space $\mathbb{CP}^{n}$ (see [BD2012, Lemma 2.3]). A triangulation of complex projective space is then obtained by triangulating the branched covering $(\mathbb{S}^{2})^{n}\to(\mathbb{S}^{2})^{n}/\operatorname{Sym}(n)=\mathbb{CP}^{n}$.
This approach is explored in [BD2011, BD2012]: Starting with the 4-vertex triangulation $S^{2}_{4}$ of $\mathbb{S}^{2}$, a simplicial subdivision $((S^{2}_{4})^{n})^{\prime}$ of the cell complex $(S^{2}_{4})^{n}$ is considered such that the quotient $((S^{2}_{4})^{n})^{\prime}/\operatorname{Sym}(n)$ is a triangulation of $\mathbb{CP}^{n}$. Carrying out this method yields a $10$-vertex triangulation of $\mathbb{CP}^{2}$ and a $30$-vertex triangulation of $\mathbb{CP}^{3}$. The latter triangulation can be reduced to an $18$-vertex triangulation of $\mathbb{CP}^{3}$ by using bistellar flips (specifically, the GAP program BISTELLAR due to Lutz [GAP2022, L2003], or any other implementation of simplification heuristics in mathematical software, such as [R2023, ES2023]). Note that, for $n\geq 3$, it is not possible to simplicially subdivide $(S^{2}_{4})^{n}$ without adding extra vertices such that the quotient is a triangulation of a manifold, cf. [BD2012].
In [KS2023], the authors follow an alternative approach to find triangulations of projective spaces: in the complex case, $\mathbb{CP}^{n}$ is decomposed into $(n+1)$ balls of real dimension $2n$ as “zones of influence”, where a given homogeneous coordinate direction is dominant in absolute value. Sets of these balls naturally intersect in points that have the corresponding set of coordinates jointly being dominant. The intersection of all balls yields a central $n$-dimensional torus. This decomposition of $\mathbb{CP}^{n}$ into $12^{n+1}-$ 1 subsets is known as its equilibrium decomposition. This point of view gives one way to triangulate $\mathbb{CP}^{n}$, by first starting with a triangulation of the $n$-torus, and then “filling in” more and more bits and pieces. This approach leads to a beautiful $10$-vertex triangulation of $\mathbb{CP}^{2}$, first described in [KB1992], but faces problems already for $n=3$.
Other approaches to construct triangulations of $\mathbb{CP}^{n}$, including a sketch of the approach presented here, are discussed in [MO2010]. Moreover, there exists an overview over triangulations of complex projective spaces in low dimensions due to Sergeraert [S2010], where the notion of triangulation denotes a more general type of decomposition with non-bijective gluing maps.
While triangulations of $\mathbb{CP}^{n}$ are notoriously difficult to explicitly write down, this is easier in the real case of $\mathbb{RP}^{n}$. Here, the first family of explicit triangulations is due to Kühnel [K1986]. The fundamental challenge that remains in the real case, is to find triangulations with as few vertices as possible. The classical triangulations from [K1986] have $12^{n+1}-$ 1 vertices, but recent advances include triangulations with $O\left(\left(\frac{1+\sqrt{5}}{2}\right)^{n+1}\right)$ vertices [VZ2021] and even a family with number of vertices subexponential in $n$ [AAK2022]. The question of whether there exist triangulations of $\mathbb{RP}^{n}$ with number of vertices a polynomial in $n$ is still open. Moreover, leaving the realm of simplicial complexes and considering the more general class of (regular) simplicial cell complexes, a very elegant and simple construction exists producing such complexes for all $\mathbb{RP}^{n}$ with the smallest number of vertices, see Example 3.8.
In this article, we follow exactly this strategy for triangulations of $\mathbb{CP}^{n}$. Namely, we again exploit the action of $\operatorname{Sym}(n)$ on $(\mathbb{S}^{2})^{n}$, but apply it to (regular) simplicial cell decompositions of $(\mathbb{S}^{2})^{n}$ and $\mathbb{CP}^{n}$ (rather than simplicial complexes): Let $S^{2}_{3}$ be the simplicial cell complex consisting of two triangles, glued along their boundaries. Taking the $n$-fold Cartesian product of $S^{2}_{3}$ with itself, we obtain a cell decomposition $\widetilde{X}^{n}$ of $(\mathbb{S}^{2})^{n}$ with $2^{n}$ cells of real dimension $2n$. We introduce the notion of a good action of a group $G\leq\operatorname{Aut}(X)$ on a simplicial cell complex $X$, and prove that if such an action is good, then $X/G$ is also a simplicial cell complex and $|X/G|$ is homeomorphic to $|X|/G$ (cf. Lemma 3.5 and Corollary 3.7). We then construct a simplicial subdivision $X^{n}$ (Theorem 4.1) of $\widetilde{X}^{n}$ without adding new vertices such that the action of $\operatorname{Sym}(n)$ on $X^{n}$ is good (cf. Theorem 5.3). The result is a simplicial cell structure $X^{n}/\operatorname{Sym}(n)$ of $\mathbb{CP}^{n}$ for all $n$ (cf. Theorem 5.4). In fact, it can be shown that $X^{n}$ can be represented as an edge coloured version of its dual graph, and thus is a graph encoded manifold, see [FGG1986] for details. Taking the first derived subdivision of the simplicial cell complex $X^{n}/\operatorname{Sym}(n)$ one obtains a triangulation of $\mathbb{CP}^{n}$ for each $n\geq 2$. This yields, for the first time, an explicit triangulation of $\mathbb{CP}^{n}$ for each $n\geq 4$.
Code to produce simplicial cell complexes $X^{n}$ and $X^{n}/\operatorname{Sym}(n)$ is discussed in Sections 6 and A, and given in [DS2024]. In Appendix B, we list $X^{n}/\operatorname{Sym}(n)$, $n\in\{2,3,4\}$ explicitly as Regina [R2023] triangulations and/or graph encodings.
Acknowledgements
Research of the second author is supported in part under the Australian Research Council’s Discovery funding scheme (project number DP220102588). This article was completed while the second author was on sabbatical at the Technische Universität (TU) Berlin. The authors would like to thank TU Berlin, and particularly Michael Joswig and his research group, for their hospitality in times of difficult circumstances.
2 Preliminaries
When we say $X$ is a poset, we mean $X$ is a finite set with a partial ordering $\leq$. By $\alpha<\beta$, we mean $\alpha\leq\beta$ and $\alpha\neq\beta$. A bijection $f:(K,\leq)\to(L,\leq)$ between two posets is called an isomorphism if $a<b$ implies $f(a)<f(b)$ for $a,b\in K$. Two posets are called isomorphic if there exits an isomorphism between them. In this article, we do not distinguish between isomorphic posets. An isomorphism $f:(K,\leq)\to(K,\leq)$ is called an automorphism of $(K,\leq)$. The set $\operatorname{Aut}(K)$ of all the automorphisms of $(K,\leq)$ is a group under composition and is called the automorphism group of $(K,\leq)$.
A CW-complex is said to be regular if all of its closed cells are homeomorphic to closed balls. Given a CW-complex $M$, let ${\mathcal{M}}$ be the set of all closed cells of $M$ together with the empty set $\emptyset$. Then ${\mathcal{M}}$ is a poset, where the partial ordering is given by set inclusion. This poset ${\mathcal{M}}$ is said to be the face poset of $M$. If $M$ and $N$ are two finite regular CW-complexes with isomorphic face posets then $M$ and $N$ are homeomorphic. A regular CW-complex $M$ is said to be simplicial, if the boundary of each cell of $M$ is isomorphic (as a poset) to the boundary of a simplex. A simplicial CW-complex is also called a simplicial cell complex. All CW-complexes considered in this paper are finite.
A simplicial poset $K=(K,\leq)$ of dimension $d$ is a poset isomorphic to the face poset ${\mathcal{M}}$ of a $d$-dimensional simplicial cell complex $M$. The topological space $M$ is called the geometric carrier of $K$ and is also denoted by $|K|$. If a topological space $N$ is homeomorphic to $|K|$, then $K$ is said to be a simplicial cell decomposition (or pseudotriangulation) of $N$. If $|K|$ is connected then we say $K$ is connected. Simplicial posets are also referred to in the literature as CW posets, pseudo-simplicial complexes or pseudocomplexes (see [BD2014, B1984, M2013] for details). We identify a simplicial cell complex $M$ with the face poset of $M$. Accordingly, by a simplicial cell complex we also mean a simplicial poset.
For $\alpha,\beta\in K$, we say $\alpha$ is incident to $\beta$ if $\alpha\leq\beta$ or $\beta\leq\alpha$. For $\alpha\in K$, $\partial\alpha:=\{\beta\in K\,:\,\beta<\alpha\}$ is a simplicial poset and is said to be the boundary of $\alpha$. If $\partial\alpha$ is isomorphic to the boundary complex of a $j$-simplex, then $\alpha$ is a called a $j$-cell of $K$. The 0-cells and 1-cells of a simplicial poset $K$ are called the vertices and edges of $K$ respectively. The set $V(K)$ of vertices of a simplicial poset $K$ is called the vertex set of $K$. The vertex set $V_{K}(\alpha)$ (:={$:=\{v\in V(K):v\leq\alpha\}$) of a $j$-cell $\alpha$ has $j+1$ vertices. If $U\subseteq V_{K}(\alpha)$, then the (unique) face of $\alpha$ whose vertex set is $U$ is denoted by $\alpha(U)$.
For an element $\alpha$ of a simplicial poset $K$, the poset $\operatorname{Lk}_{K}(\alpha):=\{\beta\in K\,:\,\alpha\leq\beta\}$ is also a simplicial poset and is called the link of $\alpha$ in $K$ (cf. [M2013]). Observe that $\alpha$ is the smallest element (i.e., element of dimension $-1$) in $\operatorname{Lk}_{K}(\alpha)$. Note that for this definition to work it is essential that the cell complex underlying $K$ is regular.
To gain some geometric intuition about the link, suppose $K$ is the face poset of a simplicial CW-complex $M$. We can assume that $M$ is a metric space. For each vertex $u$ of $K$, consider a sphere $S_{u}$ in $M$ with centre $u$ and of small radius so that $S_{u}$ intersects only members of $\operatorname{Lk}_{K}(u)\setminus\{u\}$ and $u$ is the only vertex interior of $S_{u}$. Then $L_{u}:=\{\alpha\cap S_{u}\,:\,\alpha\in M\}\cup\{\emptyset\}$ is a simplicial CW-complex, and for each $j$-cell $\beta$ of $M$ with $u\leq\beta$, $S_{u}\cap\beta$ is a $(j-1)$-cell of $L_{u}$. This shows that $\operatorname{Lk}_{K}(u)$ is isomorphic to the face poset of $L_{u}$. Now suppose that $\alpha$ is a $j$-cell, $j\geq 1$. Let $u$ be a vertex of $\alpha$. Then $\alpha$ is a $(j-1)$-cell in $\operatorname{Lk}_{K}(u)$ and $\operatorname{Lk}_{K}(\alpha)=\operatorname{Lk}_{\operatorname{Lk}_{K}(u)}(\alpha)$. By induction on the dimension of the cell $\alpha$, $\operatorname{Lk}_{\operatorname{Lk}_{K}(u)}(\alpha)$ is a simplicial cell complex and hence $\operatorname{Lk}_{K}(\alpha)$ is so.
Suppose $L\subseteq K$ are two simplicial posets. We say that $L$ is a sub-poset of $K$ if $\beta$ is a $j$-cell of $K$ and $\beta\in L$ implies the dimension of $\beta$ in $L$ is also $j$. Equivalently, if $K$ is the face poset of a simplicial CW-complex $M$ then $L$ is the face poset of a subcomplex $N$ of $M$. Thus, for $\alpha\in K$, $\partial\alpha$ is a sub-poset of $K$ but $\operatorname{Lk}_{K}(\alpha)$ is not for $\alpha\neq\emptyset$.
If $K$ is a $d$-dimensional simplicial poset and $k\leq d$ then $\operatorname{skel}_{k}(K):=\{\alpha\in K\,:\,\dim(\alpha)\leq k\}$ is a sub-poset of $K$ (with induced partial ordering) of dimension $k$ and is called the $k$-skeleton of $K$. The $1$-skeleton of $K$ is a graph and is called the edge graph of $K$. Clearly, $K$ is connected if and only if the graph $\operatorname{skel}_{1}(K)$ is connected. We say two vertices are adjacent in $K$ if they are in $\operatorname{skel}_{1}(K)$. If a vertex $v$ is a vertex in $k$ edges of $K$ then we say $v$ is a vertex of degree $k$ and we write $\deg_{K}(v)=k$.
A $d$-dimensional simplicial poset $K$ is called pure if each cell of $K$ is a face of a $d$-cell (i.e., all the maximal cells are $d$-cells). The $d$-cells and $(d-1)$-cells of a pure $d$-dimensional simplicial poset $K$ are called the facets and ridges of $K$ respectively. If $K$ is pure and $\alpha\in K$ then $\operatorname{Lk}_{K}(\alpha)$ is also pure.
The dual graph $\Lambda(K)$ of a pure simplicial poset $K$ is the graph whose vertices are the facets of $K$ and edges are the ordered pairs ($(\{\sigma_{1},\sigma_{2}\},\gamma)$, where $\gamma$ is a ridge and is a common face of the facets $\sigma_{1}$, $\sigma_{2}$. Observe that $\Lambda(K)$ is in general a multigraph without loops.
Suppose $(X,\leq)$ is the face poset of the simplicial CW-complex $M$. Let $M^{\prime}$ be the derived subdivision of $M$ (with the cells of $M$ identified with the vertices of $M^{\prime}$). The face poset of $M^{\prime}$ is called the derived subdivision of $(X,\leq)$ and is denoted by $(X^{\prime},\leq)$. The poset $X^{\prime}$ is (isomorphic to) the set of all chains of the form $\emptyset=\alpha_{0}<\alpha_{1}<\cdots<\alpha_{k}$, for some $k\geq 0$, in $(X,\leq)$ and $C_{1}\leq C_{2}$ if $C_{1}$ is a sub-chain of $C_{2}$. Since $M^{\prime}$ is a geometric simplicial complex, it follows that $(X^{\prime},\leq)$ is a simplicial complex. In fact, if $\beta_{1},\dots,\beta_{m}$ are common elements of two chains $C,D$ and $\beta_{1}<\cdots<\beta_{m}$ then $E:=\emptyset<\beta_{1}<\cdots<\beta_{m}$ is a common sub-chain of $C,D$ and is in $X^{\prime}$.
For a $d$-dimensional simplicial poset $K$, let $f_{j}(K)$ denote the number of $j$-cells of $K$ for $0\leq j\leq d$. The number $\chi(K):=\sum_{j=0}^{d}(-1)^{j}f_{j}(K)$ is called the Euler characteristic of $K$. The $(d+1)$-tuple $f(K):=(f_{0}(K),\dots,f_{d}(K))$ is called the face vector or $f$-vector of $K$.
3 Suitability for a simplicial cell complex to allow quotients
In this section we discuss the effect of taking the quotient of a simplicial cell complex by a group of automorphisms acting on it.
When considering a group $G$ acting on a simplicial complex $K$, the notions of goodness and pureness are known. Each one implies that $|K|/G\cong\left|K/G\right|$ (cf. [BD2012]). Where, the latter expression for the quotient is defined as the simplicial complex with vertex set the $G$-orbits of $V(K)$. More precisely, we have the following.
Definition 3.1.
Let $G$ be a group of automorphisms of a simplicial complex $K$. The action of $G$ on $K$ is called good if, for every pair $x\neq y$ of vertices of $K$ from a common $G$-orbit,
  • $xy$ is a non-edge; and
  • there exists a $g\in G$ such that $g(x)=y$, and $g$ fixes all the vertices of $\operatorname{Lk}(x)\cap\operatorname{Lk}(y)$.
Proposition 3.2 (Corollary 2.7 in [BD2012]).
Let $G$ be a group of automorphisms of a simplicial complex $K$. If the action of $G$ on $K$ is good, then $|K/G|\cong|K|/G$.
We require a version of Proposition 3.2 for simplicial cell complexes. For this we define
Definition 3.3.
Let $G\leq\operatorname{Aut}(X)$ be a group of automorphisms of a simplicial cell complex $X$. We say that the action of $G$ on $X$ is good, if two vertices of $X$ in the same $G$-orbit are not adjacent in $X$.
In the next lemma, we present an equivalent form (which we need in the proof of 3.6) of this customised definition of goodness.
Lemma 3.4.
Let $G\leq\operatorname{Aut}(X)$ be a group of automorphisms of a simplicial cell complex $X$. The following are equivalent.
  1. $(i)$
    For $\sigma\in X$ and $g\in G$, if $\alpha$ is a cell in $g(\sigma)\cap\sigma$, then $g(\alpha)=\alpha$.
    [[[[Observe that $g(\sigma)\in X$, but $\sigma\cap g(\sigma)$ need not be a single cell of $X$.]]]]
  2. $(ii)$
    The action of $G$ on $X$ is good.
Proof.
First assume $(i)$ holds. If possible, suppose $(ii)$ does not hold. Then there exist two vertices $x\neq y$ in the same $G$-orbit which are adjacent via an edge $e$ of $X$. Since $x$ and $y$ are in the same $G$-orbit, there exists $g\in G$ such that $g(x)=y$. Then $y$ is a 0-cell in $e\cap g(e)$. Therefore, by $(i)$, $g(y)=y$. Thus, $g(x)=y=g(y)$. This is not possible, since $g$ is an automorphism. Thus $(ii)$ holds.
Now, suppose $(ii)$ holds. Let $\sigma\in X$ and $g\in G$. Suppose $v$ is a vertex in $\sigma\cap g(\sigma)$. Then $v$ is of $\sigma$ and $v=g(u)$ for some vertex $u$ of $\sigma$. If possible, suppose $u\neq v$. Then $u,v\in V(\sigma)$ and hence vertices of the edge $e=\sigma(\{u,v\})\leq\sigma$. Hence, $v=g(u)$ and $u$ are in the same $G$-orbit and vertices of the edge $e$. This is not possible since $(ii)$ holds. Therefore, $v=u=g^{-1}(v)$, and hence $g(v)=v$. Thus, $g(x)=x$ for each vertex $x$ in $\sigma\cap g(\sigma)$.
Now suppose $\alpha$ is a cell in $\sigma\cap g(\sigma)$. Then $\alpha\leq\sigma$ and $\alpha=g(\beta)$ for some $\beta\leq\sigma$. Let $x$ be a vertex of $\alpha\leq\sigma$. Then $g^{-1}(x)$ is a vertex of $g^{-1}(\alpha)=\beta\leq\sigma$. This implies that $x$ is a vertex of $g(\sigma)$. Thus, $x$ is a vertex in $\sigma\cap g(\sigma)$. Therefore, by the above, $g(x)=x$. Hence $x=g^{-1}(x)\in V(g^{-1}(\alpha))=V(\beta)$. This implies that $V(\alpha)=V(\beta)$. Therefore, $\alpha=\sigma(V(\alpha))=\sigma(V(\beta))=\beta$. Thus, $(i)$ holds. ∎
If $G\leq\operatorname{Aut}(Y)$ for a poset $(Y,\leq)$, then consider the set $Y/G$ of $G$-orbits of elements of $Y$. The partial order $\leq$ on $Y$ induces a canonical partial order on $Y/G$ (which we also denoted by $\leq$). Namely, we have $A\leq B$, for $A,B\in Y/G$, if there exist $a\in A$ and $b\in B$ such that $a\leq b$. If $a_{0}<\cdots<a_{m}$ is a chain in $Y$, then $O_{a_{0}}<\cdots<O_{a_{m}}$ is a chain in $Y/G$. This implies that $\dim(Y/G)=\dim(Y)$. For a simplicial cell complex $(Y,\leq)$ and $G\leq\operatorname{Aut}(Y)$, $(Y/G,\leq)$ need not be a simplicial cell complex. Here we prove.
Lemma 3.5.
Let $X$ be a simplicial cell complex, and let $G\leq\operatorname{Aut}(X)$. The quotient poset $X/G$ is a simplicial cell complex if and only if the action of $G$ on $X$ is good.
Proof.
First assume that the action of $G$ on $X$ is good. We may assume that $X$ is a simplicial CW-complex. It follows that $X/G$ is a CW-complex as well. Let $\rho:X\to X/G$ be the canonical projection. Let $A\in X/G$. Then $A=\rho(\alpha)$ for some simplex $\alpha\in X$. If $a,b$ are two vertices of $\alpha$, then $a,b$ are vertices of the edge $e:=\alpha(\{a,b\})\leq\alpha$. Since the action of $G$ on $X$ is good, it follows that orbits of $a$ and $b$ are distinct. This means $\rho(a)\neq\rho(b)$ for each pair of vertices $a,b$ of $\alpha$. Therefore, $A=\rho(\alpha)$ is a cell whose boundary is isomorphic to the boundary of $\alpha$. This implies that $X/G$ is a simplicial cell complex and the dimension of $X/G$ is the same as that of $X$.
Conversely, suppose that the action of $G$ on $X$ is not good. Then there exist vertices $x,y$ of $X$ in the same $G$-orbit which are vertices of an edge $e$ of $X$. This implies that the 1-cell $\rho(e)$ is a loop in $X/G$. Therefore $X/G$ not regular and hence not a simplicial cell complex. This completes the proof. ∎
For a simplicial cell complex $X$, let $X^{\prime}$ be its derived subdivision. So, if $\alpha\in V(X^{\prime})$, then $\alpha\in X$ and $V(\operatorname{Lk}_{X^{\prime}}(\alpha))=\{\beta\in X\,:\,\beta<\alpha\}\sqcup\{\gamma\in X\,:\,\alpha\leq\gamma\}$. The group $\operatorname{Aut}(X)$ acts canonically on $X^{\prime}$ (namely, for $\{\alpha_{1}<\cdots<\alpha_{k}\}\in X^{\prime}$ and $g\in G$, $g(\{\alpha_{1}<\cdots<\alpha_{k}\}):=\{g(\alpha_{1})<\cdots<g(\alpha_{k})\}$). Hence, we can assume that $\operatorname{Aut}(X)\leq\operatorname{Aut}(X^{\prime})$. In general, $\operatorname{Aut}(X)\neq\operatorname{Aut}(X^{\prime})$. For example, if $S^{1}_{2}$ is the simplicial cell complex with two vertices and two edges, then $\operatorname{Aut}(S^{1}_{2})\cong\mathbb{Z}_{2}\times\mathbb{Z}_{2}$. Whereas $(S^{1}_{2})^{\prime}$ is a 4-cycle whose automorphism group is the dihedral group $D_{8}$ with eight elements.
Lemma 3.6.
Let $X$ be a simplicial cell complex with first derived subdivision $X^{\prime}$, and let $G\leq\operatorname{Aut}(X)$. If the action of $G$ on $X$ is good, then the action of $G$ on the simplicial complex $X^{\prime}$ is also good.
Proof.
Suppose two vertices $\alpha$ and $\beta$ of $X^{\prime}$ are in the same $G$-orbit. Let $\beta=g(\alpha)$, where $g\in G$. Then $\alpha$ and $\beta$ are cells of $X$ of the same dimension. Hence, $\alpha\not<\beta$ and $\beta\not<\alpha$ as cells of $X$. This implies that $\alpha\beta$ is a non-edge of $X^{\prime}$.
Claim 1: There does not exist $\gamma\in X$ such that $\alpha\leq\gamma$ and $\beta\leq\gamma$.
If possible, suppose there exists $\gamma\in X$ such that $\alpha\leq\gamma$, $\beta\leq\gamma$. Since a subset $A\subseteq V(\gamma)$ uniquely determines the face $\gamma(A)$ of $\gamma$ whose vertex set is $A$, it follows that $V(\alpha)\neq V(\beta)$. Since $g(\alpha)=\beta$, we have $g(V(\alpha))=V(\beta)$ and hence #$\#(V(\alpha))=\#(V(\beta))$. These imply that there exists $x\in V(\alpha)\setminus V(\beta)$. Then $g(x)\in V(g(\alpha))=V(\beta)$, and hence $g(x)\not=x$. Now, $x\leq\alpha\leq\gamma$ and $g(x)\leq\beta\leq\gamma$. Thus, $g(x)$ and $x$ are two distinct vertices of $\gamma$ and hence are vertices of the edge $\gamma(\{x,g(x)\})$. This is not possible since the action of $G$ on $X$ is good. This proves Claim 1.
Now, suppose $\nu\in V(\textrm{lk}_{X^{\prime}}(\alpha))\cap V(\textrm{lk}_{X^{\prime}}(\beta))$. By Claim 1 above, at least one of $\alpha$ and $\beta$ is not a face of $\nu$. Assume, without loss, that $\alpha\not\leq\nu$. Since $\nu\in V(\operatorname{Lk}_{X^{\prime}}(\alpha))$, it follows from the definition of the derived subdivision that $\nu<\alpha$ and hence #$\#(V(\nu))<\#(V(\alpha))$. Since $\alpha$ and $\beta$ are cells of the same dimension, we have #$\#(V(\alpha))=\#(V(\beta))$ and hence #$\#(V(\nu))<\#(V(\beta))$. This implies that $\beta\not\leq\nu$. Again, since $\nu\in V(\operatorname{Lk}_{X^{\prime}}(\beta))$, it follows that $\nu<\beta$. Thus, $\nu$ is a cell in $\alpha\cap\beta=\alpha\cap g(\alpha)$. Since the action of $G$ on $X$ is good, by Lemma 3.4, it follows that $g(\nu)=\nu$. Thus, $g$ fixes all the vertices in $\textrm{lk}_{X^{\prime}}(\alpha)\cap\textrm{lk}_{X^{\prime}}(\beta)$. Hence, the action of $G$ on the simplicial complex $X^{\prime}$ is good. ∎
We have the following consequence of Lemma 3.6.
Corollary 3.7.
Let $X$ be a simplicial cell complex, and let $G\leq\operatorname{Aut}(X)$. If the action of $G$ on $X$ is good, then $\left|X/G\right|\cong|X|/G$.
Proof.
Since $X$ is a simplicial cell complex, $X^{\prime}$ is a simplicial complex and hence $X^{\prime}/G$ is a simplicial complex. Now, by Lemma 3.5, $X/G$ is a simplicial cell complex and hence $(X/G)^{\prime}$ is a simplicial complex.
Since $G\leq\operatorname{Aut}(X)$ we have $G\leq\operatorname{Aut}(X^{\prime})$. Observe that $V(X^{\prime}/G)$ is the set of vertices of $X^{\prime}/G$ and hence the set of $G$-orbits of vertices of $X^{\prime}$. This in particular implies that $V(X^{\prime}/G)$ is the set of $G$-orbits of cells of $X$, and thus equal to $V((X/G)^{\prime})$. Moreover, the identity map from $V(X^{\prime}/G)$ to $V((X/G)^{\prime})$ gives an isomorphism from $X^{\prime}/G$ to $(X/G)^{\prime}$ in the following way: $A\in X^{\prime}/G$Longrightarrow $A=\{O_{\alpha_{1}},\dots,O_{\alpha_{k}}\}$ for some cells $\alpha_{1},\dots,\alpha_{k}\in X$, where $\alpha_{1}<\cdots<\alpha_{k}$Longrightarrow $O_{\alpha_{1}},\dots,O_{\alpha_{k}}\in X/G=V((X/G)^{\prime})$ and $O_{\alpha_{1}}<\cdots<O_{\alpha_{k}}$Longrightarrow $\{O_{\alpha_{1}},\dots,O_{\alpha_{k}}\}\in(X/G)^{\prime}$Longrightarrow $A\in(X/G)^{\prime}$. Similarly, $A\in(X/G)^{\prime}$Longrightarrow $A\in X^{\prime}/G$. Thus, $X^{\prime}/G\cong(X/G)^{\prime}$ and hence $|X^{\prime}/G|\cong|(X/G)^{\prime}|$.
Now, by Lemma 3.6, the action of $G$ on $X^{\prime}$ is good. Hence, by Proposition 3.2, $|X^{\prime}|/G\cong|X^{\prime}/G|$. Therefore, $|X|/G=|X^{\prime}|/G\cong|X^{\prime}/G|\cong|(X/G)^{\prime}|=|X/G|$. This completes the proof. ∎
We end this section with the following beautiful example of good action of a group on a very well known simplicial cell complex.
Example 3.8.
For $n\geq 1$, let $\rho:\mathbb{S}^{n}\to\mathbb{S}^{n}$ be the antipodal map. Then $\mathbb{S}^{n}/\langle\rho\rangle$ is the real projective space $\mathbb{RP}^{n}$. Let $K=S^{0}_{2}(\{u_{0},v_{0}\})\ast\cdots\ast S^{0}_{2}(\{u_{n},v_{n}\})$ be the join of $n+1$ copies of $S^{0}_{2}$. (This $K$ is known as the boundary of the $(n+1)$-dimensional cross polytope.) Then $K$ is a triangulation of $\mathbb{S}^{n}$. The map $\rho$ induces an order two automorphism $\alpha$ of $K$ given by $\alpha|_{V(K)}=(u_{0},v_{0})\cdots(u_{n},v_{n})$. More explicitly, for a simplex $x_{i_{0}}\cdots x_{i_{k}}\in K$ of dimension $k\leq n$, $\alpha(x_{i_{0}}\cdots x_{i_{k}})=y_{i_{0}}\cdots y_{i_{k}}$, where $\{x_{i_{j}},y_{i_{j}}\}=\{u_{i_{j}},v_{i_{j}}\}$ for all $j$. Let $G:=\langle\alpha\rangle\leq\operatorname{Aut}(K)$. Then, in the category of simplicial complexes, the action of $G$ on the simplicial complex $K$ is not good and as a simplicial complex $K/G$ is (the closure of) an $n$-simplex.
In the category of simplicial cell complexes, however, the action of $G$ on the simplicial cell complex $K$ is good and $K/G$ is a $(n+1)$-vertex simplicial cell structure of $\mathbb{RP}^{n}$. In particular, this produces a crystallization (simplicial cell structure with the smallest possible number of vertices) of $\mathbb{RP}^{n}$ for all $n>0$.
Remark 3.9.
Lemma 3.5 shows that the good actions on the category of simplicial cell complexes are natural actions.
4 Simplicial cell decompositions of $(\mathbb{S}^{2})^{n}$
In this section we present a very natural simplicial cell decomposition of the $n$-fold Cartesian power $(\mathbb{S}^{2})^{n}$ of the $2$-sphere $\mathbb{S}^{2}$.
For this, consider the following well-known method to simplicially subdivide the $n$-fold Cartesian product of a triangle: Let the triangle be $\Delta=x_{1}x_{2}x_{3}$. Then the vertices of the cell $\Delta^{n}$ are $x_{i_{1}\cdots i_{n}}:=(x_{i_{1}},\dots,x_{i_{n}})$, $i_{j}\in\{1,2,3\}$. Take a subdivision of an $n$-cube $C^{n}=[1,3]^{n}$ into $2^{n}$ sub-cubes. Let $C^{\prime}_{n}$ be the union of all these $2^{n}$ cubes, i.e., $C_{n}^{\prime}=([1,2]\cup[2,3])^{n}$. The vertices of $C^{\prime}_{n}$ are $(i_{1},i_{2},\dots,i_{n})$, where $i_{j}\in\{1,2,3\}$ for $1\leq j\leq n$. We identify $(i_{1},i_{2},\dots,i_{n})$ with $x_{i_{1}i_{2}\cdots i_{n}}$. A (maximal) monotone path from $x_{11\cdots 1}=(1,1,\dots,1)$ to $x_{33\cdots 3}=(3,3,\dots,3)$ in the edge graph of $C^{\prime}_{n}$ is a path of the form $x_{11\cdots 1}\mbox{-}\cdots\mbox{-}x_{i_{1}i_{2}\cdots i_{n}}\mbox{-}x_{j_{1}j_{2}\cdots j_{n}}\mbox{-}\cdots\mbox{-}x_{33\cdots 3}$, where $i_{k}\leq j_{k}$ for all $k$, and $j_{1}+\cdots+j_{n}=i_{1}+\cdots+i_{n}+1$. The facets of the simplicial subdivision of $\Delta^{n}$ are given by the vertex sets of maximal monotone paths. Such a path contains $2n$ edges and hence $12n+$ 1 vertices, and hence these facets are, indeed, of dimension $2n$ (see Figure 1 for an illustration of the case $n=2$; see [BCS1988, Page 382] and [ES1952, Page 67, Definition 8.8] for more details). To calculate the number of facets in this simplicial subdivision of $\Delta^{n}$, observe that, to reach $(3,\ldots,3)$ from $(1,\ldots,1)$, we must follow every dimension of the cube exactly twice. There are (${2n\choose 2}$ possibilities to decide when to go along the first dimension, (${2n-2\choose 2}$ options for the second, etc. It follows that we have $\prod\limits_{j=0}^{n-1}\binom{2(n-j)}{2}=\frac{(2n)!}{2^{n}}$ facets in the subdivision.
Figure 1: The dual graph of the subdivision of $\Delta^{2}\times\Delta^{2}$ into $6$ simplices of dimension $4$. Notice that this subdivision is a graph encoding.
Let $S^{2}_{3}$ be the standard (unique) $3$-vertex $2$-triangle crystallisation of $\mathbb{S}^{2}$ (a pair of triangles, identified along their boundaries). We refer to the two triangles of $S^{2}_{3}$ as $U$ (for upper) and $L$ (for lower). Taking the $n$-fold topological Cartesian product of $S^{2}_{3}$ with itself, we obtain a cell decomposition $\widetilde{X}^{n}$ of $(\mathbb{S}^{2})^{n}$ with $2^{n}$ facets. Each facet is an $n$-fold Cartesian product of triangles and is uniquely described by a binary word whose letters are in $\{U,L\}$. We identify the facet $A_{1}\times\cdots\times A_{n}$ with the word $A_{1}\cdots A_{n}$, where $A_{1},\dots,A_{n}\in\{U,L\}$. We have for the boundary of $\partial(A_{1}\cdots A_{n})=\cup_{k=1}^{n}(A_{1}\times\cdots\times A_{k-1}\times(\partial A_{k})\times A_{k+1}\times\cdots\times A_{n})$. Thus, a facet has exactly $3n$ boundary ridges, each of them a Cartesian product of $n-1$ triangles and one edge. Moreover, the facets $A_{1}\ldots A_{n}$ and $B_{1}\cdots B_{n}$ share exactly three ridges if and only if $A_{i}=B_{i}$ for all but exactly one index $j=1,\ldots,n$, and $A_{j}\neq B_{j}$. Otherwise they share none. These three ridges are exactly the ones determined by the $3$ edges of the $j$-th triangle.
We simplicially subdivide each cell of the complex $\widetilde{X}^{n}$ into ${(2n)!}/{2^{n}}$ facets, using the monotone path method given above. If carried out in all $2^{n}$ maximal cells, we obtain a total of $2^{n}\times((2n)!/2^{n})=(2n)!$ facets for the subdivision of $\widetilde{X}^{n}$. The subdivision is consistent on the ridges of $\widetilde{X}^{n}$ for the following reason: Let the vertices of $S^{2}_{3}$ be $x_{1},x_{2},x_{3}$. Moreover, let $F=A_{1}\cdots A_{n}$ be a maximal cell of $\widetilde{X}^{n}$, and let $F^{\prime}$ be its subdivision into simplices. Naturally, $F$ contains all $3^{n}$ vertices of $\widetilde{X}^{n}$. Each of its ridges $F_{\ell;i,j}=A_{1}\times\cdots\times A_{\ell-1}\times x_{i}x_{j}\times A_{\ell+1}\times\cdots\times A_{n}$, $1\leq i<j\leq 3$ and $1\leq\ell\leq n$ is incident to exactly one other maximal cell $E=A_{1}\cdots A_{\ell-1}B_{\ell}A_{\ell+1}\cdots A_{n}$, where $\{A_{\ell},B_{\ell}\}=\{U,L\}$. By construction, we have $F_{\ell;i,j}=E_{\ell;i,j}$, and the restriction of the simplicial subdivision $F^{\prime}$ of $F$ on $F_{\ell;i,j}$ is equal to the restriction of the simplicial subdivision $E^{\prime}$ of $E$ on $F_{\ell;i,j}$. To see this, note that the $(2n-1)$-simplices in $F_{\ell;i,j}$ are the maximal monotone paths in the edge graph of the subdivision $C^{n}_{\ell;i,j}=([1,2]\cup[2,3])^{\ell-1}\times[i,j]\times([1,2]\cup[2,3])^{n-\ell}$ of the $n$-dimensional rectangular parallelepiped (cube if $(i,j)=(1,3)$) $[1,3]^{\ell-1}\times[i,j]\times[1,3]^{n-\ell}$.
Altogether this yields the following.
Theorem 4.1.
The simplicial cell complex
[…]
$\mbox{of }C^{\prime}_{n},~{}V(P)\mbox{ is the vertex set of }P,\mbox{ and }F\mbox{ a maximal cell of }\widetilde{X}^{n}\}$
is a simplicial cell decomposition of $({\mathbb{S}}^{2})^{n}$.
Example 4.2.
We have the following $f$-vectors of $X^{n}$, for $n\in\{1,2,3,4\}$:
  • $f(X^{1})=(3,3,2)$,
  • $f(X^{2})=(9,27,58,60,24)$,
  • $f(X^{3})=(27,189,926,2460,3504,2520,720)$,
  • $f(X^{4})=(81,1215,12130,64860,194280,337680,338400,181440,40320)$.
In general, $X^{n}$ has $3^{n}$ vertices, $(2n)!$ facets, and Euler characteristic $2^{n}$. Moreover, $X^{n}$ can be represented as a graph encoding. That is, as an edge-coloured dual graph containing all the information to recover the simplicial cell complex. This can be followed from the construction, but we do not particularly stress this in this article. See Appendix A for code to produce these simplicial cell complexes in low-dimensional topology software Regina [R2023].
5 Simplicial cell decompositions of $\mathbb{CP}^{\hskip 0.85358ptn}$
In this section we use the simplicial cell decomposition $X^{n}$ of $(\mathbb{S}^{2})^{n}$ from Section 4 to construct a simplicial cell decomposition of $\mathbb{CP}^{n}$. For this, we first define an action of the group $\operatorname{Sym}(n)$ on $X^{n}$ realising coordinate permutation, and then show that this action is good. The quotient $X^{n}/\operatorname{Sym}(n)$ is our required simplicial cell complex.
The natural action of the symmetric group $\operatorname{Sym}(n)$ of degree $n$ by coordinate permutation on $(\mathbb{S}^{2})^{n}$ is also an action on $(S^{2}_{3})^{n}$ by coordinate permutation. This gives an action of $\operatorname{Sym}(n)$ on $\widetilde{X}^{n}$: if $\pi\in\operatorname{Sym}(n)$ and $A=\alpha_{1}\times\cdots\times\alpha_{n}$ is a cell (not necessarily facet) in $\widetilde{X}^{n}$ then $\pi(\alpha_{1}\times\cdots\times\alpha_{n})=\alpha_{\pi(1)}\times\cdots\times\alpha_{\pi(n)}$.
Let $C_{n}=[1,3]^{n}$ and $C_{n}^{\prime}=([1,2]\cup[2,3])^{n}$ be as in Section 4. In particular, the vertices of $C^{\prime}_{n}$ are $x_{i_{1}i_{2}\cdots i_{n}}:=(i_{1},i_{2},\dots,i_{n})$, $i_{j}\in\{1,2,3\}$, for $1\leq j\leq n$. Then, $\operatorname{Sym}(n)$ acts on the vertices of $C_{n}^{\prime}$ (which are also the vertices of $\widetilde{X}^{n}$ and $X^{n}$) as $\pi(i_{1},\dots,i_{n})=(\ell_{1},\dots,\ell_{n})$, where $\ell_{k}=i_{\pi(k)}$, $\pi\in\operatorname{Sym}(n)$.
Here we prove
Lemma 5.1.
The action of $\operatorname{Sym}(n)$ on $V(X^{n})$ induces an action of $\operatorname{Sym}(n)$ on the simplicial cell complex $X^{n}$.
Proof.
Let $H$ be the edge graph of $C^{\prime}_{n}$. Then the distance $d_{H}(x_{i_{1}i_{2}\cdots i_{n}},x_{11\cdots 1})$ between $x_{i_{1}i_{2}\cdots i_{n}}$ and $x_{11\cdots 1}$ in $H$ is $(i_{1}-1)+\cdots+(i_{n}-1)=(i_{1}+\cdots+i_{n})-n$.
Claim 2. If $x_{i_{1}i_{2}\cdots i_{n}}$ is a vertex of $C^{\prime}_{n}$ and $\pi\in\operatorname{Sym}(n)$ then $d_{H}(x_{i_{1}i_{2}\cdots i_{n}},x_{11\cdots 1})=d_{H}(\pi(x_{i_{1}i_{2}\cdots i_{n}}),x_{11\cdots 1})$.
If $\rho=(r,s)\in\operatorname{Sym}(n)$, $r<s$, is a transposition then $\rho(x_{i_{1}\cdots i_{r}\cdots i_{s}\cdots i_{n}})=x_{i_{1}\cdots i_{r-1}i_{s}i_{r+1}\cdots i_{s-1}i_{r}i_{s+1}\cdots i_{n}}$. This implies $d_{H}(x_{i_{1}i_{2}\cdots i_{n}},x_{11\cdots 1})=d_{H}(\rho(x_{i_{1}i_{2}\cdots i_{n}}),$ $x_{11\cdots 1})$. Since any permutation is the product of transpositions, this implies that $d_{H}(x_{i_{1}i_{2}\cdots i_{n}},x_{11\cdots 1})=d_{H}(\pi(x_{i_{1}i_{2}\cdots i_{n}}),x_{11\cdots 1})$. This proves Claim 2.
Claim 3. If $uv$ is an edge of $C^{\prime}_{n}$ and $\pi\in\operatorname{Sym}(n)$ then $\pi(u)\pi(v)$ is also an edge of $C^{\prime}_{n}$.
Suppose $u=x_{i_{1}i_{2}\cdots i_{n}}$ and $v=x_{j_{1}j_{2}\cdots j_{n}}$. Since $uv$ is an edge, there exists $\ell\in\{1,\dots,n\}$ such that $i_{k}=j_{k}$ for all $k\in\{1,\dots,n\}\setminus\{\ell\}$ and $|i_{\ell}-j_{\ell}|=1$. Assume without loss that $j_{\ell}=i_{\ell}+1$. Hence, $j_{1}+\cdots+j_{n}=i_{1}+\cdots+i_{n}+1$. Then $i_{\pi(k)}=j_{\pi(k)}$ for all $k\in\{1,\dots,n\}\setminus\{\ell\}$. Since $\sum_{r=1}^{n}i_{r}=\sum_{r=1}^{n}i_{\pi(r)}$ and $\sum_{r=1}^{n}j_{r}=\sum_{r=1}^{n}j_{\pi(r)}$, these imply $j_{\pi(\ell)}=\sum_{r=1}^{n}j_{r}-(j_{\pi(1)}+\cdots+j_{\pi(\ell-1)}+j_{\pi(\ell+1)}+\cdots+j_{\pi(n)})=(\sum_{r=1}^{n}i_{r})+1-(i_{\pi(1)}+\cdots+i_{\pi(\ell-1)}+i_{\pi(\ell+1)}+\cdots+i_{\pi(n)})=1+i_{\pi(\ell)}$. These imply $\pi(u)\pi(v)=x_{i_{\pi(1)}\cdots i_{\pi(n)}}x_{j_{\pi(1)}\cdots j_{\pi(n)}}$ is an edge of $C^{\prime}_{n}$. This proves Claim 3.
Let $P=v_{0}\mbox{-}v_{1}\mbox{-}\cdots\mbox{-}v_{2n-1}\mbox{-}v_{2n}$ be a maximal monotone path and $\pi\in\operatorname{Sym}(n)$. So, $v_{0}=x_{1\cdots 1}=\pi(x_{1\cdots 1})=\pi(v_{0})$ and $v_{2m}=x_{3\cdots 3}=\pi(x_{3\cdots 3})=\pi(v_{2m})$. Then, by Claim 3, $\pi(v_{i-1})\pi(v_{i})$ is an edge of $C^{\prime}_{n}$ and by Claim 2, $d_{H}(\pi(v_{i}),\pi(v_{0}))=i$ for $1\leq i\leq 2m$. This implies that $\pi(v_{0})\mbox{-}\pi(v_{1})\mbox{-}\cdots\mbox{-}\pi(v_{2n-1})\mbox{-}\pi(v_{2n})$ is a maximal monotone path in the edge graph of $C^{\prime}_{n}$. Therefore, for any maximal monotone path $P$ and $\pi\in\operatorname{Sym}(n)$, $\pi(V(P))$ is the vertex set of another maximal monotone path. Therefore, for any facet $V(P)_{F}$ of $X^{n}$, $\pi(V(P))_{\pi(F)}$ is a facet of $X^{n}$. These imply that $\pi(A_{F})=\pi(A)_{\pi(F)}$, for $A_{F}\in X^{n}$ and $\pi\in\operatorname{Sym}(n)$, gives an action of $\operatorname{Sym}(n)$ on $X^{n}$. This proves the lemma. ∎
The proof of Lemma 5.1 implies the following statements.
Definition 5.2.
The action of $\operatorname{Sym}(n)$ on $\widetilde{X}^{n}$ induces the following action on $X^{n}$. If $\{v_{1},\dots,v_{m}\}_{F}$ is a simplex of $X^{n}$ and $\pi\in\operatorname{Sym}(n)$ then
$\pi(\{v_{1},\dots,v_{m}\}_{F})=\{\pi(v_{1}),\dots,\pi(v_{m})\}_{\pi(F)}.$
We are now in a position to state and prove our main results.
Theorem 5.3.
The action of   $\operatorname{Sym}(n)$ on $X^{n}$ given in Definition 5.2 is good.
Proof.
Claim 4. Let $x_{11\cdots 1}\mbox{-}\cdots\mbox{-}x_{i_{1}i_{2}\cdots i_{n}}\mbox{-}\cdots\mbox{-}x_{k_{1}k_{2}\cdots k_{n}}\mbox{-}\cdots\mbox{-}x_{33\cdots 3}$ be a monotone path. Then there does not exist $\pi\in\operatorname{Sym}(n)$ such that $\pi(x_{i_{1}i_{2}\cdots i_{n}})=x_{k_{1}k_{2}\cdots k_{n}}$.
Since $x_{11\cdots 1}\mbox{-}\cdots\mbox{-}x_{i_{1}i_{2}\cdots i_{n}}\mbox{-}\cdots\mbox{-}x_{k_{1}k_{2}\cdots k_{n}}\mbox{-}\cdots\mbox{-}x_{33\cdots 3}$ is a monotone path, it follows that $d_{H}(x_{i_{1}i_{2}\cdots i_{n}},x_{11\cdots 1})<d_{H}(x_{k_{1}k_{2}\cdots k_{n}},x_{11\cdots 1})$. Now, if $\pi\in\operatorname{Sym}(n)$ then, by Claim 2 above, $d_{H}(\pi(x_{i_{1}i_{2}\cdots i_{n}}),x_{11\cdots 1})=d_{H}(x_{i_{1}i_{2}\cdots i_{n}},x_{11\cdots 1})<d_{H}(x_{k_{1}k_{2}\cdots k_{n}},$ $x_{11\cdots 1})$. Hence, $\pi(x_{i_{1}i_{2}\cdots i_{n}})\neq x_{k_{1}k_{2}\cdots k_{n}}$. This proves Claim 4.
Suppose the action of $\operatorname{Sym}(n)$ on $X^{n}$ is not good. Then there exists a pair of vertices $x,y$ in a $\operatorname{Sym}(n)$-orbit which are vertices of an edge $e$. Since each edge is incident to a facet, we have a facet $\sigma$ such that $e\leq\sigma$. This implies, $x,y$ are vertices of $\sigma$ (in $F^{\prime}$ for some maximal cell $F$ of $\widetilde{X}^{n}$). Thus, $x,y$ are vertices of a maximal monotone path and are in a $\operatorname{Sym}(n)$-orbit. But this is not possible by Claim 4. This proves the result. ∎
Now we have
Theorem 5.4.
The quotient $X^{n}/\operatorname{Sym}(n)$ is a simplicial cell complex and $|X^{n}/\operatorname{Sym}(n)|\cong\mathbb{CP}^{n}$ for all $n\geq 2$.
Proof.
By Theorem 5.3 and Lemma 3.5, $X^{n}/\operatorname{Sym}(n)$ is a simplicial cell complex. This proves the first part.
From Theorem 5.3 and Corollary 3.7, we get $|X^{n}/\operatorname{Sym}(n)|\cong|X^{n}|/\operatorname{Sym}(n)$. Therefore, $|X^{n}/\operatorname{Sym}(n)|\cong|X^{n}|/\operatorname{Sym}(n)\cong|\widetilde{X}^{n}|/\operatorname{Sym}(n)\cong(\mathbb{S}^{2})^{n}/\operatorname{Sym}(n)\cong\mathbb{CP}^{n}$. This proves the result. ∎
Moreover, we have the following consequence of Theorem 5.4.
Corollary 5.5.
The simplicial complex $(X^{n})^{\prime}/\operatorname{Sym}(n)$ is a triangulation of $\mathbb{CP}^{n}$ for all $n\geq 2$.
Proof.
First note that $\operatorname{Sym}(n)\leq\operatorname{Aut}(X^{n})\leq\operatorname{Aut}((X^{n})^{\prime})$. Now, from the proof of Corollary 3.7, $(X^{n})^{\prime}/\operatorname{Sym}(n)$ and $(X^{n}/\operatorname{Sym}(n))^{\prime}$ are isomorphic simplicial complexes. The result now follows from Theorem 5.4. ∎
Example 5.6.
We have the following $f$-vectors of $T_{n}=X^{n}/\operatorname{Sym}(n)$, for $n\in\{1,2,3,4\}$:
  • $f(T_{1})=(3,3,2)$,
  • $f(T_{2})=(6,15,30,30,12)$,
  • $f(T_{3})=(10,46,184,440,596,420,120)$,
  • $f(T_{4})=(15,111,764,3345,8982,14700,14280,7560,1680)$.
In general, $T_{n}$ has (${n+2\choose 2}$ vertices, $(2n)!/n!$ facets, and Euler characteristic $n+1$. As in the case of $X^{n}$ (see Example 4.2), $T_{n}$ can be represented as a graph encoding. That is, as an edge-coloured dual graph containing all the information to recover the simplicial cell complex. See Appendix A for code to produce these simplicial cell complexes in low-dimensional topology software Regina [R2023].
6 Implementation
One advantage of our construction of simplicial cell decompositions of $(\mathbb{S}^{2})^{n}$ and $\mathbb{CP}^{n}$ is that they can be explicitly implemented. In Appendix A, we provide python3 [P2009] code to produce simplicial cell decompositions for $n\in\{2,3,4\}$ within low-dimensional topology software Regina [R2023] [^1] In Regina, what we call simplicial cell decomposition is called a triangulation. We sometimes switch between the two terms whenever it is clear what we mean from context.. This code is also available from [DS2024]. Since Regina’s python interface only implements simplicial cell decompositions up to real dimension eight, we added code to produce a graph encoding of our simplicial cell decompositions which works for arbitrary values of $n$.
Moreover, in Appendix B we provide our simplicial cell structures of $\mathbb{CP}^{2}$, $\mathbb{CP}^{3}$, and $\mathbb{CP}^{4}$ explicitly in the form of Regina (pseudo) triangulations or graph encodings (or both).
Appendix A Code to produce $X^{n}$ and $X^{n}/\operatorname{Sym}(n)$
A.1 Helper functions to produce our simplicial cell decompositions
This code has been tested within the python interface of Regina regina-python. It is also available from [DS2024]. First, we need the help of the following two python packages.
# for permutation iterators
import itertools
# for computing the inverse of a permutation
from sympy.combinatorics import Permutation
Then we need to a function to produce all path orderings for our simplicial subdivision of the $n$-fold cartesian product of the triangle.
def orderings(all_orderings,pos):
␣␣Thisfunctionproducesallmonotonepathsdefiningthesimplicesin
␣␣thecanonicaltriangulationofthen-foldcartesianproductofatriangle
␣␣Foragivennthefunctionmustbeinitiatedwith
␣␣all_orderings=listoflength2nfilledwith’-1
␣␣pos=1
␣␣outputisalistoflistsoflength2nfilledwithentries1,...,n
␣␣lengthoflistis(2nchoose2)x(2n-2choose2)x(4choose2)
  next_level = []
  # go through all partial paths found so far, and fill in the next value (pos)
  for lst in all_orderings:
    for i in range(len(lst)):
      # skip entries that are already filled
      if lst[i] != -1:
        continue
      # skip entries that are already filled
      for j in range(i+1,len(lst)):
        if lst[j] != -1:
          continue
        # fill every pair of unfilled entries with value ’pos’
        tmp = lst.copy()
        tmp[i] = pos
        tmp[j] = pos
        # add to new list of partial paths
        next_level.append(tmp)
  if not -1 in next_level[0]:
    # if done, return complete list of paths in cube
    return next_level
  else:
    # otherwise, call next level of recursion
    return orderings(next_level,3*pos)
With this, the simplicial subdivision of the $n$-fold product of the triangle is obtained by the following function (which produces an abstract simplicial complex)
def triangle_product(n):
␣␣Thisfunctionproducesallsimplicesinthecanonicaltriangulationof
␣␣then-foldCartesianproductofthetrianglewithvertices0,....,3^n-1
␣␣numberoffacetsisthesameasnumberofmonotonepathsreturnedby
␣␣orderings(.,.)((2nchoose2)x(2n-2choose2)x(4choose2)=2n!/2^n)
  facets = []
  ords = orderings([[-1]*(2*n)],1)
  ords.sort()
  # turn paths into simplicial complex subdividing the n-fold Cartesian product of the triangle
  for ordering in ords:
    f=[0]
    for k in ordering:
      f.append(f[-1]+k)
    facets.append(f)
  return facets
The following function produces a list of boundary faces of the subdivided triangle product.
def boundary_faces(facets):
␣␣Thisfunctionreturnsthe(Z2)-boundaryofasimplicialcomplexdefinedby
␣␣"facets"
  # record co-dim. one boundary faces of simplices as we see them
  faces = []
  # record their position as a pair [facet,boundary face index]
  pos = []
  for f in range(len(facets)):
    for k in range(len(facets[f])):
      tmp = facets[f].copy()
      tmp.remove(facets[f][k])
      if not tmp in faces:
        # add new boundary faces
        faces.append(tmp)
        pos.append([f,k])
      else:
        # remove boundary faces if we encounter them a second time
        pos.remove(pos[faces.index(tmp)])
        faces.remove(tmp)
  return faces,pos
This abstract simplicial complex needs to be converted into a Regina (pseudo) triangulation.
def simpToReg(simp,d,triangulation,perm):
␣␣TurnfacetlistofsimplicialcomplexintoReginatriangulation
␣␣Here,thisisonlyappliedtothesimplicialsubdivisionofthen-foldCartesian
␣␣productofatriangle(2xn=d)
␣␣simp␣␣␣␣␣␣␣␣␣␣=facetlist(ofabstractsimplicialcomplex)
␣␣d␣␣␣␣␣␣␣␣␣␣␣␣␣=dimensionoftriangulation(d=len(simp[0])-1)
␣␣triangulation=appropriatereginafunctionforgeneratingad-dimensionaltriangulation
␣␣perm␣␣␣␣␣␣␣␣␣␣=appropriatereginafunctionforpermutingelementsofasimplex
␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣(d+1)points
  facets = []
  for i in range(d+1):
    tmp = list(range(d+1))
    tmp.remove(i)
    facets.append(tmp)
  n = len(simp)
  neighbours = []
  gluings = []
  for j in range(n):
    neighbours.append([None] * (d+1))
    gluings.append([None] * (d+1))
    for f in range((d+1)):
      found = False
      for k in range(n):
        if k == j:
          continue
        good = True
        for i in range(d):
          if simp[j][facets[f][i]] not in simp[k]:
            good = False
            break
        if not good:
          continue
        found = True
        break
      if not found:
        continue
      neighbours[j][f] = k
      tmp = [None] * (d+1)
      unused = int(d*(d+1)/2)
      for i in range(d):
        adj = simp[k].index(simp[j][facets[f][i]])
        tmp[facets[f][i]] = adj
        unused = unused - adj
      tmp[f] = unused
      gluings[j][f] = perm(tmp)
  ans = triangulation()
  for j in range(n):
      ans.newSimplex()
  for j in range(n):
    for f in range(d+1):
      if ans.simplex(j).adjacentSimplex(f) == None and gluings[j][f] != None:
        ans.simplex(j).join(f,ans.simplex(neighbours[j][f]), gluings[j][f])
  return ans
And finally we need code to produce the orbits of simplices of length $n!$ given by the $\operatorname{Sym}(n)$-action that is (complex) coordinate permutation
def orbits(n):
␣␣Thisfunctionreturnsthesetoforbits(eachoflengthn!)ofthenatural
␣␣Sym(n)-actiononthecoordinatesofthen-foldCartesianproductof
␣␣the2-triangle2-sphere
␣␣Ideaisthatanorbitismadeoutofallpermutationsofthelettersof
␣␣thedefiningorbit(onthelevelofasimplexinacell),with
␣␣targetcelldefinedbypermutingthelettersofthecelllabel
␣␣Forthis,weusethefollowingglobalcountingofsimplices:
␣␣1.triangulationofonecellhasorderingofsimplicesiaccording
␣␣␣␣␣tolexicographicorderingofpathlabels(Nsimplicesoverall)
␣␣2.ForcellB_j={U,L}^n,wecannumberfromallU=0toallL=2^n-1
␣␣3.Globalindexofsimplexisthenidx=i+Nxj
␣␣Orbits:
␣␣␣(A)Everyorbitmustcontainonerepresentativeofeverypermutationofpaths
␣␣␣(B)Thisgivesgroupsoforbitsofn!pathsin{nchoose#Us}={nchoose#Ls}
␣␣␣␣␣␣␣facets
␣␣␣(C)Ifasimplexisgivenbyapathpinacellc,then,forallsigmain
␣␣␣␣␣␣␣Sym(n):producesigma(pxc)
␣␣Coordinateaxesareencodedasfollows:3^i~x_(i+1)
  lut = [1]
  for k in range(1,n):
    lut.append(3*lut[-1])
  lut.reverse()
  # list for orbits (each entry of length n!, (2n)!/n! orbits)
  orbs = []
  # paths defining simplices and number of simplices in triangulation of cell
  paths = orderings([[-1]*(2*n)],1)
  paths.sort()
  cell_len = len(paths) # this is equal to int(math.factorial(2*n)/2**n)
  # for conversion of cell number to binary
  frmt = "0"+str(n)+"b"
  # symmetric group on n elements
  all_perms = list(itertools.permutations(list(range(n))))
  # loop over all cells c
  for c in range(2**n):
    # loop over all simplices in cell c
    for q in range(cell_len):
      # check if we have seen this facet already
      already_done = False
      for o in orbs:
        if cell_len*c + q in o:
          already_done = True
          break
      if already_done:
        continue
      # found a representative for an orbit not yet processed
      # convert cell number to binary
      bnr = format(c, frmt)
      # collect group of orbits (orbit in cell) x (orbit of cells)
      all_bnrs = list(set(list(itertools.permutations(list(bnr)))))
      all_bnrs.sort()
      for seed in all_bnrs:
        cur_orb = []
        # permute paths by all permutation on n letters
        for p in range(len(all_perms)):
          # variable for permuted path
          new_path = []
          for k in paths[q]:
            new_path.append(lut[all_perms[p][lut.index(k)]])
          # determine target cell (permute cell)
          new_seed = list(seed)
          # take inverse of permutation
          sigma = Permutation(list(all_perms[p]))**(-1)
          # binary word for cell is in reverse with respect to coordinate axes enumeration
          new_bnr = []
          new_seed.reverse()
          for k in list(sigma):
            new_bnr.append(new_seed[k])
          new_bnr.reverse()
          # represent cell label as binary string for conversion
          P = ’’.join(new_bnr)
          dec = int(P,2)
          # add new facet in target cell to orbit
          cur_orb.append(dec*cell_len+paths.index(new_path))
        cur_orb.sort()
        orbs.append(cur_orb)
  orbs.sort()
  return orbs
A.2 Code to produce sphere products
With the helper functions defined in Section A.1 in place, we can now present code to produce our simplicial cell decompositions of $(\mathbb{S}^{2})^{n}$, $n\in\{2,3,4\}$.
def sphere_product(NN,triangulation,perm):
␣␣returnsphereproduct(worksuptoNN=4,reginaonlyallowstriangulations
␣␣upto8=2x4realdimensions)
␣␣NN␣␣␣␣␣␣␣␣␣␣␣␣=(complex)dimensionofcomplexprojectivespace
␣␣triangulation=appropriatereginafunctionforgeneratingad-dimensionaltriangulation
␣␣perm␣␣␣␣␣␣␣␣␣␣=appropriatereginafunctionforpermutingelementsofasimplex
␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣(d+1)points
  O = orderings([[-1]*(2*NN)],1)
  O.sort()
  F = triangle_product(NN)
  faces,pos = boundary_faces(F)
  len_faces = len(F)
  Tri = simpToReg(F,2*NN,triangulation,perm)
  # for conversion to binary
  frmt = "0"+str(NN)+"b"
  # build up triangulation from disjoint triangulations of cells
  # cells are numbered by which number they represent in binary - 1
  T = triangulation()
  for c in range(2**NN):
    T.insertTriangulation(Tri)
  # loop over cells
  for c in range(2**NN):
    # make buckets of free faces to connect to neighbour cells:
    buckets = []
    for i in range(NN):
      buckets.append([])
    for j in range(len(faces)):
      for char in range(NN-1):
        # Algorithm:
        # char=0: len=2: first
        # char=0, len=3: not first (done for n=2)
        #  char=1, len=4: second
        #  char=1, len=5: not first not second (done for n=3)
        #    char=2, len=6: third
        #    char=2, len=7: not first, second, third (done for n=4)
        # for entry 2*(char+1)
        ctr1 = 0
        # for entry 2*(char+1)+1
        ctr2 = 0
        res=[]
        for k in faces[j]:
          if not k%3**(char+1) in res:
            res.append(k%3**(char+1))
        if len(res) == 2*(char+1):
          buckets[char].append(pos[j])
          break
        if len(res) == 2*(char+1)+1 and char == NN-2:
          buckets[char+1].append(pos[j])
    # decide to which cell to glue to
    bnr = format(c, frmt)
    for char in range(NN):
      # already glued
      if bnr[char] == ’1’:
        continue
      ll = list(bnr)
      ll[char] = ’1’
      newBnr=’’.join(ll)
      newC = int(newBnr,2)
      for buc in buckets[char]:
        T.simplex(len_faces*c+buc[0]).join(buc[1],T.simplex(buc[0]+len_faces*newC),perm())
  return T
The sphere products in dimensions $4$, $6$, and $8$ can then be called as follows.
def s2xs2():
␣␣returnsatriangulationofS2xS2
  return sphere_product(2,Triangulation4,Perm5)
def s2xs2xs2():
␣␣returnsatriangulationofS2xS2xS2
  return sphere_product(3,Triangulation6,Perm7)
def s2xs2xs2xs2():
␣␣returnsatriangulationofS2xS2xS2xS2
  return sphere_product(4,Triangulation8,Perm9)
A.3 Code to produce $\mathbb{CP}^{n}$
Finally, we can produce the code to construct simplicial cell decompositions $T_{n}=X^{n}/\operatorname{Sym}(n)$, for $n\in\{2,3,4\}$:
def CP(NN,triangulation,perm):
  # calculate orbits
  orbs = orbits(NN)
  # compute sphere product
  sp = sphere_product(NN,triangulation,perm)
  # empty triangulation to contain complex projective space
  T = triangulation()
  # compute paths for simplicial subdivision of single cell
  paths = orderings([[-1]*(2*NN)],1)
  # sort (IMPORTANT to do this after every call of ’orderings(.,.)’)
  paths.sort()
  # go through orbits and add one simplex per orbit
  for o in range(len(orbs)):
    T.newSimplex()
  # for every smallest orbit representative: detect gluing from sphere product sp
  for k in range(len(orbs)):
    # go through all orbit reps and check their adjacent orbits
    for simp in orbs[k]:
      for l in range(2*NN+1):
        adj = sp.simplex(simp).adjacentSimplex(l)
        target = False
        for o in range(len(orbs)):
          if adj.index() in orbs[o]:
            target = o
            break
    rep = orbs[k][0]
    for l in range(2*NN+1):
      adj = sp.simplex(rep).adjacentSimplex(l)
      # if face is still free
      if T.simplex(k).adjacentSimplex(l) == None:
        # get target simplex
        for o in range(len(orbs)):
          if adj.index() in orbs[o]:
            target = o
            break
        if T.simplex(target).adjacentSimplex(l) == None:
          T.simplex(k).join(l,T.simplex(target),perm())
  return T
The complex spaces in (real) dimensions $4$, $6$, and $8$ can then be called as follows.
def CP2():
␣␣returnsatriangulationofCP2
  return CP(2,Triangulation4,Perm5)
def CP3():
␣␣returnsatriangulationofCP3
  return CP(3,Triangulation6,Perm7)
def CP4():
␣␣returnsatriangulationofCP4
  return CP(4,Triangulation8,Perm9)
Appendix B Explicit presentation of our simplicial cell complexes
In this section we provide automorphism signatures of our (pseudo) triangulations of $(\mathbb{S}^{2})^{n}$, and $\mathbb{CP}^{n}$, $n\in\{1,2,3,4\}$ which can be used to directly construct the (pseudo) triangulations using Regina [R2023] command T = TriangulationN.fromIsoSig() (where N is to be replaced by the appropriate real dimension).
B.1 $\mathbb{S}^{2}\times\mathbb{S}^{2}$ and $\mathbb{CP}^{2}$
Isomorphism signature of $\mathbb{S}^{2}\times\mathbb{S}^{2}$:
yLwMMLvAzQQvQAQAPzQMQbceeefgjjkkmnonlpqrqppsqsruuuwvwvxxxxaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaa
aaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaa
Triangulation of $\mathbb{S}^{2}\times\mathbb{S}^{2}$ as a graph encoding:
Colour 0:
((0,12),(1,13),(2,14),(3,9),(4,10),(5,11),(6,18),(7,19),(8,20),(15,21),(16,22),(17,23))
Colour 1:
((0,12),(1,3),(2,4),(5,11),(6,18),(7,9),(8,10),(13,15),(14,16),(17,23),(19,21),(20,22))
Colour 2:
((0,1),(2,8),(3,15),(4,5),(6,7),(9,21),(10,11),(12,13),(14,20),(16,17),(18,19),(22,23))
Colour 3:
((0,6),(1,2),(3,4),(5,17),(7,8),(9,10),(11,23),(12,18),(13,14),(15,16),(19,20),(21,22))
Colour 4:
((0,6),(1,7),(2,14),(3,9),(4,16),(5,17),(8,20),(10,22),(11,23),(12,18),(13,19),(15,21))
Isomorphism signature of $\mathbb{CP}^{2}$:
mLwMMLLAQQQbceeefgkjjlkjklljlkaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaa
Simplicial cell decomposition of $\mathbb{CP}^{2}$ as a graph encoding:
Colour 0:
((0,8),(1,7),(2,6),(3,9),(4,10),(5,11))
Colour 1:
((0,8),(1,2),(3,9),(4,6),(5,7),(10,11))
Colour 2:
((0,1),(2,5),(3,4),(6,11),(7,8),(9,10))
Colour 3:
((0,3),(1,2),(4,5),(6,7),(8,9),(10,11))
Colour 4:
((0,3),(1,4),(2,6),(5,11),(7,10),(8,9))
B.2 $\mathbb{S}^{2}\times\mathbb{S}^{2}\times\mathbb{S}^{2}$ and $\mathbb{CP}^{3}$
The isomorphism signature of $(S^{2})^{3}$ is somewhat too long to present it here. It can be generated using the code in [DS2024]. The isomorphism signature of $\mathbb{CP}^{3}$:
-c4bLLMLzLPLQwvALAPAvMQQQwQMQAvvPwAALMQMAQPAQAPQQQQAPQPwzzAQLwQMAQQPAPAQQMQLQMPQLQQMQAvwA
MQQQAAAAQLQwQQPQPPQQQQQPAvMAQAMAQQQQPQPQAQQwPMQQPQQQAQQcaeadafahafajajajahakalaoanapapaqa
ravasasawaxaqazawaBaraDaCawauaAaAatayaEaBauaGaFazaCaGaGaLaJaNaKaOaHaHaSaIaVaTaMaJaVaNaKaK
aOaPaWaQaQaUaXaZaRaYaYaHaQa0aIaYaWa1aRaWaVaLaRaKaKaIa0aSaXaZa1aTa3a3aYaVa5a6a9a9a+a-a-a+a
8a-a-acbcbabdbdbab9acb8a7aebgbdb8a5afbbb6aibcb7a7afbbbib8adb9aebfbmb8agbbbmb5akbebgb6akbo
bhbhbqblbmbibqbrbnbpbjbrbnbtbubvbtbvbzbxbzbxbxbubzbxbzbvbybxbxbzbubBbubybvbBbsbFbDbybFbAb
wbGbAbBbtbHbIbCbDbDbybJbKbEbAbKbJbsbsbIbCbGbEbwbHbEbJbFbFbKbGbLbKbNbObPbPbQbRbQbRbMbPbRbS
bPbTbUbObObUbNbObMbSbRbTbMbSbVbSbTbQbTbUbNbNbWbMbMbSbXbWbXbVbVbXbZbZbZb1b0b1b1b1bYb1bZbZb
2b0bYbYb2b0b0b2b2bYbYb3b3b3b3b3b3baaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaa
aaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaa
aaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaa
aaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaa
aaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaa
aaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaa
aaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaa
aaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaa
aaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaa
aaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaa
aaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaaa
The simplicial cell decomposition of $\mathbb{CP}^{3}$ represented as a graph encoding has two automorphisms of order $2$. The first leaves colours invariant, the second permutes colour $i$ to $6-i$, $0\leq i\leq 6$.
The automorphisms are given as
(1,106)(2,107)(3,108)(4,109)(5,110)(6,111)(7,112)(8,113)(9,114)(10,115)(11,116)(12,117)
(13,118)(14,119)(15,120)(16,93)(17,92)(18,91)(19,90)(20,89)(21,88)(22,102)(23,101)(24,100)
(25,105)(26,104)(27,103)(28,97)(29,96)(30,99)(31,98)(32,94)(33,95)(34,81)(35,80)(36,79)
(37,85)(38,84)(39,87)(40,86)(41,82)(42,83)(43,76)(44,77)(45,78)(46,69)(47,68)(48,67)
(49,73)(50,72)(51,75)(52,74)(53,70)(54,71)(55,64)(56,65)(57,66)(58,61)(59,62)(60,63)
for the colour preserving one, and
(2,4)(3,7)(6,8)(11,12)(16,58)(17,55)(18,43)(19,48)(20,36)(22,59)(23,56)(24,44)(25,60)
(26,57)(27,45)(28,53)(29,41)(30,54)(31,42)(34,47)(37,50)(39,52)(61,93)(62,102)(63,105)
(64,92)(65,101)(66,104)(67,90)(68,81)(70,97)(71,99)(72,85)(74,87)(76,91)(77,100)(78,103)
(79,89)(82,96)(83,98)(107,109)(108,112)(111,113)(116,117)
for the one shuffling colours. Orbit representatives are then given for the first four colours 00 to $3$ as follows:
Colour 0:
(1,58),(2,59),(3,60),(4,55),(5,56),(6,57),(7,48),(8,47),(9,46),(10,53),(11,54),(12,50),
(13,49),(14,52),(15,51),(16,88),(17,89),(18,90),(22,79),(23,80),(24,81),(25,76),(26,77),
(27,78),(28,84),(29,85),(30,82),(31,83),(32,87),(33,86)
Colour 1:
(1,58),(2,59),(3,60),(4,7),(5,8),(6,9),(10,12),(11,13),(14,15),(16,88),(17,89),(18,90),
(22,25),(23,26),(24,27),(28,30),(29,31),(32,33),(34,46),(35,47),(36,48),(37,49),(38,50),
(39,51),(40,52),(41,53),(42,54),(43,55),(44,56),(45,57)
Colour 2:
(1,4),(2,5),(3,6),(7,43),(8,44),(9,45),(10,11),(12,14),(13,15),(16,22),(17,23),(18,24),
(19,34),(20,35),(21,36),(25,67),(26,68),(27,69),(28,37),(29,38),(30,39),(31,40),(32,41),
(33,42),(49,51),(50,52),(53,54),(55,58)(56,59),(57,60)
Colour 3:
(1,21),(2,3),(4,7),(5,10),(6,11),(8,12),(9,13),(14,32),(15,33),(16,61),(17,19),(18,20),
(22,25),(23,28),(24,29),(26,30),(27,31),(34,37),(35,38),(36,43),(39,74),(40,75),(41,44),
(42,45),(46,49),(47,50),(48,55),(53,56),(54,57),(59,60),
B.3 $\mathbb{CP}^{4}$
The simplicial cell decomposition of $(S^{2})^{4}$ has $8!=40320$ simplices of dimension eight, and is, hence, too large to be presente here. It can be generated using the code in [DS2024] either as a Regina [R2023] (pseudo) triangulation or as a gem.
The simplicial cell decomposition of $\mathbb{CP}^{4}$ has $1680$ simplices, we choose the representation as a graph encoding. Moreover, we use the fact that the simplicial cell decomposition has two automorphisms of order $2$. The first leaves colours invariant, the second permutes colour $i$ to $8-i$, $0\leq i\leq 8$.
The automorphisms are given as
(   1,1576)(   2,1577)(   3,1578)(   4,1579)(   5,1580)(   6,1581)(   7,1582)(   8,1583)
(   9,1584)(  10,1585)(  11,1586)(  12,1587)(  13,1588)(  14,1589)(  15,1590)(  16,1591)
(  17,1592)(  18,1593)(  19,1594)(  20,1595)(  21,1596)(  22,1597)(  23,1598)(  24,1599)
(  25,1600)(  26,1601)(  27,1602)(  28,1603)(  29,1604)(  30,1605)(  31,1606)(  32,1607)
(  33,1608)(  34,1609)(  35,1610)(  36,1611)(  37,1612)(  38,1613)(  39,1614)(  40,1615)
(  41,1616)(  42,1617)(  43,1618)(  44,1619)(  45,1620)(  46,1621)(  47,1622)(  48,1623)
(  49,1624)(  50,1625)(  51,1626)(  52,1627)(  53,1628)(  54,1629)(  55,1630)(  56,1631)
(  57,1632)(  58,1633)(  59,1634)(  60,1635)(  61,1636)(  62,1637)(  63,1638)(  64,1639)
(  65,1640)(  66,1641)(  67,1642)(  68,1643)(  69,1644)(  70,1645)(  71,1646)(  72,1647)
(  73,1648)(  74,1649)(  75,1650)(  76,1651)(  77,1652)(  78,1653)(  79,1654)(  80,1655)
(  81,1656)(  82,1657)(  83,1658)(  84,1659)(  85,1660)(  86,1661)(  87,1662)(  88,1663)
(  89,1664)(  90,1665)(  91,1666)(  92,1667)(  93,1668)(  94,1669)(  95,1670)(  96,1671)
(  97,1672)(  98,1673)(  99,1674)( 100,1675)( 101,1676)( 102,1677)( 103,1678)( 104,1679)
( 105,1680)( 106,1383)( 107,1382)( 108,1381)( 109,1380)( 110,1379)( 111,1378)( 112,1392)
( 113,1391)( 114,1390)( 115,1395)( 116,1394)( 117,1393)( 118,1387)( 119,1386)( 120,1389)
( 121,1388)( 122,1384)( 123,1385)( 124,1371)( 125,1370)( 126,1369)( 127,1375)( 128,1374)
( 129,1377)( 130,1376)( 131,1372)( 132,1373)( 133,1366)( 134,1367)( 135,1368)( 136,1359)
( 137,1358)( 138,1357)( 139,1363)( 140,1362)( 141,1365)( 142,1364)( 143,1360)( 144,1361)
( 145,1354)( 146,1355)( 147,1356)( 148,1351)( 149,1352)( 150,1353)( 151,1473)( 152,1472)
( 153,1471)( 154,1470)( 155,1469)( 156,1468)( 157,1482)( 158,1481)( 159,1480)( 160,1485)
( 161,1484)( 162,1483)( 163,1477)( 164,1476)( 165,1479)( 166,1478)( 167,1474)( 168,1475)
( 169,1461)( 170,1460)( 171,1459)( 172,1465)( 173,1464)( 174,1467)( 175,1466)( 176,1462)
( 177,1463)( 178,1456)( 179,1457)( 180,1458)( 181,1503)( 182,1502)( 183,1501)( 184,1500)
( 185,1499)( 186,1498)( 187,1512)( 188,1511)( 189,1510)( 190,1515)( 191,1514)( 192,1513)
( 193,1507)( 194,1506)( 195,1509)( 196,1508)( 197,1504)( 198,1505)( 199,1491)( 200,1490)
( 201,1489)( 202,1495)( 203,1494)( 204,1497)( 205,1496)( 206,1492)( 207,1493)( 208,1486)
( 209,1487)( 210,1488)( 211,1548)( 212,1547)( 213,1546)( 214,1551)( 215,1550)( 216,1549)
( 217,1543)( 218,1542)( 219,1545)( 220,1544)( 221,1540)( 222,1541)( 223,1560)( 224,1559)
( 225,1558)( 226,1563)( 227,1562)( 228,1561)( 229,1555)( 230,1554)( 231,1557)( 232,1556)
( 233,1552)( 234,1553)( 235,1572)( 236,1571)( 237,1570)( 238,1575)( 239,1574)( 240,1573)
( 241,1567)( 242,1566)( 243,1569)( 244,1568)( 245,1564)( 246,1565)( 247,1525)( 248,1524)
( 249,1527)( 250,1526)( 251,1522)( 252,1523)( 253,1531)( 254,1530)( 255,1533)( 256,1532)
( 257,1528)( 258,1529)( 259,1537)( 260,1536)( 261,1539)( 262,1538)( 263,1534)( 264,1535)
( 265,1516)( 266,1517)( 267,1518)( 268,1519)( 269,1520)( 270,1521)( 271,1413)( 272,1412)
( 273,1411)( 274,1417)( 275,1416)( 276,1419)( 277,1418)( 278,1414)( 279,1415)( 280,1408)
( 281,1409)( 282,1410)( 283,1425)( 284,1424)( 285,1423)( 286,1429)( 287,1428)( 288,1431)
( 289,1430)( 290,1426)( 291,1427)( 292,1420)( 293,1421)( 294,1422)( 295,1441)( 296,1440)
( 297,1443)( 298,1442)( 299,1438)( 300,1439)( 301,1447)( 302,1446)( 303,1449)( 304,1448)
( 305,1444)( 306,1445)( 307,1453)( 308,1452)( 309,1455)( 310,1454)( 311,1450)( 312,1451)
( 313,1432)( 314,1433)( 315,1434)( 316,1435)( 317,1436)( 318,1437)( 319,1396)( 320,1397)
( 321,1398)( 322,1399)( 323,1400)( 324,1401)( 325,1402)( 326,1403)( 327,1404)( 328,1405)
( 329,1406)( 330,1407)( 331,1284)( 332,1283)( 333,1282)( 334,1288)( 335,1287)( 336,1290)
( 337,1289)( 338,1285)( 339,1286)( 340,1279)( 341,1280)( 342,1281)( 343,1276)( 344,1277)
( 345,1278)( 346,1308)( 347,1307)( 348,1306)( 349,1312)( 350,1311)( 351,1314)( 352,1313)
( 353,1309)( 354,1310)( 355,1303)( 356,1304)( 357,1305)( 358,1320)( 359,1319)( 360,1318)
( 361,1324)( 362,1323)( 363,1326)( 364,1325)( 365,1321)( 366,1322)( 367,1315)( 368,1316)
( 369,1317)( 370,1336)( 371,1335)( 372,1338)( 373,1337)( 374,1333)( 375,1334)( 376,1342)
( 377,1341)( 378,1344)( 379,1343)( 380,1339)( 381,1340)( 382,1348)( 383,1347)( 384,1350)
( 385,1349)( 386,1345)( 387,1346)( 388,1327)( 389,1328)( 390,1329)( 391,1330)( 392,1331)
( 393,1332)( 394,1291)( 395,1292)( 396,1293)( 397,1294)( 398,1295)( 399,1296)( 400,1297)
( 401,1298)( 402,1299)( 403,1300)( 404,1301)( 405,1302)( 406,1261)( 407,1262)( 408,1263)
( 409,1264)( 410,1265)( 411,1266)( 412,1267)( 413,1268)( 414,1269)( 415,1270)( 416,1271)
( 417,1272)( 418,1273)( 419,1274)( 420,1275)( 421,1194)( 422,1193)( 423,1192)( 424,1198)
( 425,1197)( 426,1200)( 427,1199)( 428,1195)( 429,1196)( 430,1189)( 431,1190)( 432,1191)
( 433,1186)( 434,1187)( 435,1188)( 436,1218)( 437,1217)( 438,1216)( 439,1222)( 440,1221)
( 441,1224)( 442,1223)( 443,1219)( 444,1220)( 445,1213)( 446,1214)( 447,1215)( 448,1230)
( 449,1229)( 450,1228)( 451,1234)( 452,1233)( 453,1236)( 454,1235)( 455,1231)( 456,1232)
( 457,1225)( 458,1226)( 459,1227)( 460,1246)( 461,1245)( 462,1248)( 463,1247)( 464,1243)
( 465,1244)( 466,1252)( 467,1251)( 468,1254)( 469,1253)( 470,1249)( 471,1250)( 472,1258)
( 473,1257)( 474,1260)( 475,1259)( 476,1255)( 477,1256)( 478,1237)( 479,1238)( 480,1239)
( 481,1240)( 482,1241)( 483,1242)( 484,1201)( 485,1202)( 486,1203)( 487,1204)( 488,1205)
( 489,1206)( 490,1207)( 491,1208)( 492,1209)( 493,1210)( 494,1211)( 495,1212)( 496,1171)
( 497,1172)( 498,1173)( 499,1174)( 500,1175)( 501,1176)( 502,1177)( 503,1178)( 504,1179)
( 505,1180)( 506,1181)( 507,1182)( 508,1183)( 509,1184)( 510,1185)( 511,1156)( 512,1157)
( 513,1158)( 514,1159)( 515,1160)( 516,1161)( 517,1162)( 518,1163)( 519,1164)( 520,1165)
( 521,1166)( 522,1167)( 523,1168)( 524,1169)( 525,1170)( 526,1063)( 527,1064)( 528,1065)
( 529,1060)( 530,1061)( 531,1062)( 532,1053)( 533,1052)( 534,1051)( 535,1058)( 536,1059)
( 537,1055)( 538,1054)( 539,1057)( 540,1056)( 541,1048)( 542,1049)( 543,1050)( 544,1041)
( 545,1040)( 546,1039)( 547,1046)( 548,1047)( 549,1043)( 550,1042)( 551,1045)( 552,1044)
( 553,1026)( 554,1025)( 555,1024)( 556,1023)( 557,1022)( 558,1021)( 559,1037)( 560,1038)
( 561,1034)( 562,1033)( 563,1036)( 564,1035)( 565,1029)( 566,1028)( 567,1027)( 568,1032)
( 569,1031)( 570,1030)( 571,1138)( 572,1139)( 573,1140)( 574,1135)( 575,1136)( 576,1137)
( 577,1128)( 578,1127)( 579,1126)( 580,1133)( 581,1134)( 582,1130)( 583,1129)( 584,1132)
( 585,1131)( 586,1153)( 587,1154)( 588,1155)( 589,1150)( 590,1151)( 591,1152)( 592,1143)
( 593,1142)( 594,1141)( 595,1148)( 596,1149)( 597,1145)( 598,1144)( 599,1147)( 600,1146)
( 601,$1$)( 602,1112)( 603,1113)( 604,1104)( 605,1103)( 606,1102)( 607,1109)( 608,1110)
( 609,1106)( 610,1105)( 611,1108)( 612,1107)( 613,1123)( 614,1124)( 615,1125)( 616,1116)
( 617,1115)( 618,1114)( 619,1121)( 620,1122)( 621,1118)( 622,1117)( 623,1120)( 624,1119)
( 625,1074)( 626,1073)( 627,1072)( 628,1077)( 629,1076)( 630,1075)( 631,1069)( 632,1068)
( 633,1071)( 634,1070)( 635,1066)( 636,1067)( 637,1094)( 638,1095)( 639,1091)( 640,1090)
( 641,1093)( 642,1092)( 643,1100)( 644,1101)( 645,1097)( 646,1096)( 647,1099)( 648,1098)
( 649,1081)( 650,1080)( 651,1083)( 652,1082)( 653,1078)( 654,1079)( 655,1087)( 656,1086)
( 657,1089)( 658,1088)( 659,1084)( 660,1085)( 661, 958)( 662, 959)( 663, 960)( 664, 951)
( 665, 950)( 666, 949)( 667, 956)( 668, 957)( 669, 953)( 670, 952)( 671, 955)( 672, 954)
( 673, 936)( 674, 935)( 675, 934)( 676, 933)( 677, 932)( 678, 931)( 679, 947)( 680, 948)
( 681, 944)( 682, 943)( 683, 946)( 684, 945)( 685, 939)( 686, 938)( 687, 937)( 688, 942)
( 689, 941)( 690, 940)( 691,1006)( 692,1007)( 693,1008)( 694, 999)( 695, 998)( 696, 997)
( 697,1004)( 698,1005)( 699,1001)( 700,1000)( 701,1003)( 702,1002)( 703,1018)( 704,1019)
( 705,1020)( 706,1011)( 707,1010)( 708,1009)( 709,1016)( 710,1017)( 711,1013)( 712,1012)
( 713,1015)( 714,1014)( 715, 969)( 716, 968)( 717, 967)( 718, 972)( 719, 971)( 720, 970)
( 721, 964)( 722, 963)( 723, 966)( 724, 965)( 725, 961)( 726, 962)( 727, 989)( 728, 990)
( 729, 986)( 730, 985)( 731, 988)( 732, 987)( 733, 995)( 734, 996)( 735, 992)( 736, 991)
( 737, 994)( 738, 993)( 739, 976)( 740, 975)( 741, 978)( 742, 977)( 743, 973)( 744, 974)
( 745, 982)( 746, 981)( 747, 984)( 748, 983)( 749, 979)( 750, 980)( 751, 858)( 752, 857)
( 753, 856)( 754, 855)( 755, 854)( 756, 853)( 757, 867)( 758, 866)( 759, 865)( 760, 870)
( 761, 869)( 762, 868)( 763, 862)( 764, 861)( 765, 864)( 766, 863)( 767, 859)( 768, 860)
( 769, 846)( 770, 845)( 771, 844)( 772, 850)( 773, 849)( 774, 852)( 775, 851)( 776, 847)
( 777, 848)( 778, 841)( 779, 842)( 780, 843)( 781, 905)( 782, 906)( 783, 902)( 784, 901)
( 785, 904)( 786, 903)( 787, 897)( 788, 896)( 789, 895)( 790, 900)( 791, 899)( 792, 898)
( 793, 923)( 794, 924)( 795, 920)( 796, 919)( 797, 922)( 798, 921)( 799, 929)( 800, 930)
( 801, 926)( 802, 925)( 803, 928)( 804, 927)( 805, 910)( 806, 909)( 807, 912)( 808, 911)
( 809, 907)( 810, 908)( 811, 916)( 812, 915)( 813, 918)( 814, 917)( 815, 913)( 816, 914)
( 817, 876)( 818, 875)( 819, 874)( 820, 880)( 821, 879)( 822, 882)( 823, 881)( 824, 877)
( 825, 878)( 826, 871)( 827, 872)( 828, 873)( 829, 888)( 830, 887)( 831, 886)( 832, 892)
( 833, 891)( 834, 894)( 835, 893)( 836, 889)( 837, 890)( 838, 883)( 839, 884)( 840, 885)
for the colour preserving one, and
(   2,  16)(   3,  31)(   5,  19)(   6,  34)(   8,  22)(   9,  37)(  10,  46)(  11,  58)
(  12,  49)(  13,  61)(  14,  70)(  15,  73)(  18,  32)(  21,  35)(  24,  38)(  25,  47)
(  26,  59)(  27,  50)(  28,  62)(  29,  71)(  30,  74)(  40,  48)(  41,  60)(  42,  51)
(  43,  63)(  44,  72)(  45,  75)(  53,  64)(  55,  66)(  56,  76)(  57,  78)(  68,  77)
(  69,  79)(  83,  88)(  85,  90)(  86,  94)(  87,  96)(  92,  95)(  93,  97)( 101, 102)
( 106, 511)( 107, 496)( 108, 406)( 109, 433)( 110, 343)( 111, 148)( 112, 514)( 113, 499)
( 114, 409)( 115, 517)( 116, 502)( 117, 412)( 118, 484)( 119, 394)( 120, 487)( 121, 397)
( 122, 319)( 123, 322)( 124, 430)( 125, 340)( 126, 145)( 127, 445)( 128, 355)( 129, 457)
( 130, 367)( 131, 280)( 132, 292)( 134, 178)( 135, 208)( 136, 423)( 137, 333)( 139, 438)
( 140, 348)( 141, 450)( 142, 360)( 143, 273)( 144, 285)( 146, 171)( 147, 201)( 149, 156)
( 150, 186)( 151, 512)( 152, 497)( 153, 407)( 154, 434)( 155, 344)( 157, 515)( 158, 500)
( 159, 410)( 160, 518)( 161, 503)( 162, 413)( 163, 485)( 164, 395)( 165, 488)( 166, 398)
( 167, 320)( 168, 323)( 169, 431)( 170, 341)( 172, 446)( 173, 356)( 174, 458)( 175, 368)
( 176, 281)( 177, 293)( 180, 209)( 181, 513)( 182, 498)( 183, 408)( 184, 435)( 185, 345)
( 187, 516)( 188, 501)( 189, 411)( 190, 519)( 191, 504)( 192, 414)( 193, 486)( 194, 396)
( 195, 489)( 196, 399)( 197, 321)( 198, 324)( 199, 432)( 200, 342)( 202, 447)( 203, 357)
( 204, 459)( 205, 369)( 206, 282)( 207, 294)( 211, 520)( 212, 505)( 213, 415)( 214, 522)
( 215, 507)( 216, 417)( 217, 490)( 218, 400)( 219, 492)( 220, 402)( 221, 325)( 222, 327)
( 223, 521)( 224, 506)( 225, 416)( 226, 523)( 227, 508)( 228, 418)( 229, 491)( 230, 401)
( 231, 493)( 232, 403)( 233, 326)( 234, 328)( 235, 524)( 236, 509)( 237, 419)( 238, 525)
( 239, 510)( 240, 420)( 241, 494)( 242, 404)( 243, 495)( 244, 405)( 245, 329)( 246, 330)
( 247, 478)( 248, 388)( 249, 480)( 250, 390)( 251, 313)( 252, 315)( 253, 479)( 254, 389)
( 255, 481)( 256, 391)( 257, 314)( 258, 316)( 259, 482)( 260, 392)( 261, 483)( 262, 393)
( 263, 317)( 264, 318)( 266, 267)( 271, 428)( 272, 338)( 274, 443)( 275, 353)( 276, 455)
( 277, 365)( 279, 290)( 283, 429)( 284, 339)( 286, 444)( 287, 354)( 288, 456)( 289, 366)
( 295, 464)( 296, 374)( 297, 470)( 298, 380)( 300, 305)( 301, 465)( 302, 375)( 303, 471)
( 304, 381)( 307, 476)( 308, 386)( 309, 477)( 310, 387)( 331, 422)( 334, 437)( 335, 347)
( 336, 449)( 337, 359)( 346, 425)( 349, 440)( 351, 452)( 352, 362)( 358, 427)( 361, 442)
( 363, 454)( 370, 461)( 372, 467)( 373, 377)( 376, 463)( 378, 469)( 382, 473)( 384, 475)
( 424, 436)( 426, 448)( 441, 451)( 462, 466)( 526,1063)( 527,1138)( 528,1153)( 529,1060)
( 530,1135)( 531,1150)( 532,1048)( 533, 958)( 534, 778)( 535,$1$)( 536,1123)( 537,1006)
( 538, 826)( 539,1018)( 540, 838)( 541,1053)( 542,1128)( 543,1143)( 544,1041)( 545, 951)
( 546, 771)( 547,1104)( 548,1116)( 549, 999)( 550, 819)( 551,1011)( 552, 831)( 553,1026)
( 554, 936)( 555, 756)( 556, 858)( 557, 678)( 559,1074)( 560,1077)( 561, 969)( 562, 789)
( 563, 972)( 564, 792)( 565, 867)( 566, 687)( 568, 870)( 569, 690)( 571,1064)( 572,1139)
( 573,1154)( 574,1061)( 575,1136)( 576,1151)( 577,1049)( 578, 959)( 579, 779)( 580,1112)
( 581,1124)( 582,1007)( 583, 827)( 584,1019)( 585, 839)( 586,1065)( 587,1140)( 588,1155)
( 589,1062)( 590,1137)( 591,1152)( 592,1050)( 593, 960)( 594, 780)( 595,1113)( 596,1125)
( 597,1008)( 598, 828)( 599,1020)( 600, 840)( 601,1058)( 602,1133)( 603,1148)( 604,1046)
( 605, 956)( 606, 776)( 607,1109)( 608,1121)( 609,1004)( 610, 824)( 611,1016)( 612, 836)
( 613,1059)( 614,1134)( 615,1149)( 616,1047)( 617, 957)( 618, 777)( 619,1110)( 620,1122)
( 621,1005)( 622, 825)( 623,1017)( 624, 837)( 625,1037)( 626, 947)( 627, 767)( 628,1038)
( 629, 948)( 630, 768)( 631, 905)( 632, 725)( 633, 906)( 634, 726)( 637,1094)( 638,1100)
( 639, 989)( 640, 809)( 641, 995)( 642, 815)( 643,1095)( 644,1101)( 645, 990)( 646, 810)
( 647, 996)( 648, 816)( 649, 923)( 650, 743)( 651, 924)( 652, 744)( 655, 929)( 656, 749)
( 657, 930)( 658, 750)( 661,1052)( 662,1127)( 663,1142)( 664,1040)( 665, 950)( 666, 770)
( 667,1103)( 668,1115)( 669, 998)( 670, 818)( 671,1010)( 672, 830)( 673,1025)( 674, 935)
( 675, 755)( 676, 857)( 679,1073)( 680,1076)( 681, 968)( 682, 788)( 683, 971)( 684, 791)
( 685, 866)( 688, 869)( 691,1055)( 692,1130)( 693,1145)( 694,1043)( 695, 953)( 696, 773)
( 697,1106)( 698,1118)( 699,1001)( 700, 821)( 701,1013)( 702, 833)( 703,1057)( 704,1132)
( 705,1147)( 706,1045)( 707, 955)( 708, 775)( 709,1108)( 710,1120)( 711,1003)( 712, 823)
( 713,1015)( 714, 835)( 715,1034)( 716, 944)( 717, 764)( 718,1036)( 719, 946)( 720, 766)
( 721, 902)( 723, 904)( 727,1091)( 728,1097)( 729, 986)( 730, 806)( 731, 992)( 732, 812)
( 733,1093)( 734,1099)( 735, 988)( 736, 808)( 737, 994)( 738, 814)( 739, 920)( 741, 922)
( 745, 926)( 747, 928)( 751,1023)( 752, 933)( 754, 855)( 757,1029)( 758, 939)( 760,1032)
( 761, 942)( 763, 897)( 765, 900)( 769, 846)( 772, 876)( 774, 888)( 781,1069)( 782,1071)
( 783, 964)( 785, 966)( 787, 862)( 790, 864)( 793,1081)( 794,1083)( 795, 976)( 797, 978)
( 799,1087)( 800,1089)( 801, 982)( 803, 984)( 805, 910)( 807, 916)( 811, 912)( 813, 918)
( 817, 850)( 820, 880)( 822, 892)( 829, 852)( 832, 882)( 834, 894)( 841,1051)( 842,1126)
( 843,1141)( 844,1039)( 845, 949)( 847,1102)( 848,1114)( 849, 997)( 851,1009)( 853,1024)
( 854, 934)( 859,1072)( 860,1075)( 861, 967)( 863, 970)( 871,1054)( 872,1129)( 873,1144)
( 874,1042)( 875, 952)( 877,1105)( 878,1117)( 879,1000)( 881,1012)( 883,1056)( 884,1131)
( 885,1146)( 886,1044)( 887, 954)( 889,1107)( 890,1119)( 891,1002)( 893,1014)( 895,1033)
( 896, 943)( 898,1035)( 899, 945)( 907,1090)( 908,1096)( 909, 985)( 911, 991)( 913,1092)
( 914,1098)( 915, 987)( 917, 993)( 931,1022)( 937,1028)( 940,1031)( 961,1068)( 962,1070)
( 973,1080)( 974,1082)( 979,1086)( 980,1088)(1156,1383)(1157,1473)(1158,1503)(1159,1392)
(1160,1482)(1161,1512)(1162,1395)(1163,1485)(1164,1515)(1165,1548)(1166,1560)(1167,1551)
(1168,1563)(1169,1572)(1170,1575)(1171,1382)(1172,1472)(1173,1502)(1174,1391)(1175,1481)
(1176,1511)(1177,1394)(1178,1484)(1179,1514)(1180,1547)(1181,1559)(1182,1550)(1183,1562)
(1184,1571)(1185,1574)(1186,1380)(1187,1470)(1188,1500)(1189,1371)(1190,1461)(1191,1491)
(1192,1359)(1193,1284)(1195,1413)(1196,1425)(1197,1308)(1198,1218)(1199,1320)(1200,1230)
(1201,1387)(1202,1477)(1203,1507)(1204,1389)(1205,1479)(1206,1509)(1207,1543)(1208,1555)
(1209,1545)(1210,1557)(1211,1567)(1212,1569)(1213,1375)(1214,1465)(1215,1495)(1216,1363)
(1217,1288)(1219,1417)(1220,1429)(1221,1312)(1223,1324)(1224,1234)(1225,1377)(1226,1467)
(1227,1497)(1228,1365)(1229,1290)(1231,1419)(1232,1431)(1233,1314)(1235,1326)(1237,1525)
(1238,1531)(1239,1527)(1240,1533)(1241,1537)(1242,1539)(1243,1441)(1244,1447)(1245,1336)
(1247,1342)(1248,1252)(1249,1443)(1250,1449)(1251,1338)(1253,1344)(1255,1453)(1256,1455)
(1257,1348)(1259,1350)(1261,1381)(1262,1471)(1263,1501)(1264,1390)(1265,1480)(1266,1510)
(1267,1393)(1268,1483)(1269,1513)(1270,1546)(1271,1558)(1272,1549)(1273,1561)(1274,1570)
(1275,1573)(1276,1379)(1277,1469)(1278,1499)(1279,1370)(1280,1460)(1281,1490)(1282,1358)
(1285,1412)(1286,1424)(1287,1307)(1289,1319)(1291,1386)(1292,1476)(1293,1506)(1294,1388)
(1295,1478)(1296,1508)(1297,1542)(1298,1554)(1299,1544)(1300,1556)(1301,1566)(1302,1568)
(1303,1374)(1304,1464)(1305,1494)(1306,1362)(1309,1416)(1310,1428)(1313,1323)(1315,1376)
(1316,1466)(1317,1496)(1318,1364)(1321,1418)(1322,1430)(1327,1524)(1328,1530)(1329,1526)
(1330,1532)(1331,1536)(1332,1538)(1333,1440)(1334,1446)(1337,1341)(1339,1442)(1340,1448)
(1345,1452)(1346,1454)(1351,1378)(1352,1468)(1353,1498)(1354,1369)(1355,1459)(1356,1489)
(1360,1411)(1361,1423)(1367,1456)(1368,1486)(1372,1408)(1373,1420)(1384,1396)(1385,1399)
(1397,1474)(1398,1504)(1400,1475)(1401,1505)(1402,1540)(1403,1552)(1404,1541)(1405,1553)
(1406,1564)(1407,1565)(1409,1462)(1410,1492)(1415,1426)(1421,1463)(1422,1493)(1432,1522)
(1433,1528)(1434,1523)(1435,1529)(1436,1534)(1437,1535)(1439,1444)(1458,1487)(1517,1518)
(1577,1591)(1578,1606)(1580,1594)(1581,1609)(1583,1597)(1584,1612)(1585,1621)(1586,1633)
(1587,1624)(1588,1636)(1589,1645)(1590,1648)(1593,1607)(1596,1610)(1599,1613)(1600,1622)
(1601,1634)(1602,1625)(1603,1637)(1604,1646)(1605,1649)(1615,1623)(1616,1635)(1617,1626)
(1618,1638)(1619,1647)(1620,1650)(1628,1639)(1630,1641)(1631,1651)(1632,1653)(1643,1652)
(1644,1654)(1658,1663)(1660,1665)(1661,1669)(1662,1671)(1667,1670)(1668,1672)(1676,1677)
for the one shuffling colours. Orbit representatives are then given for the first five colours 00 to $4$ as follows:
Colour 0:
(1,511),(2,512),(3,513),(4,514),(5,515),(6,516),(7,517),(8,518),(9,519),(10,520),(11,521),
(12,522),(13,523),(14,524),(15,525),(16,496),(17,497),(18,498),(19,499),(20,500),(21,501),
(22,502),(23,503),(24,504),(25,505),(26,506),(27,507),(28,508),(29,509),(30,510),(31,433),
(32,434),(33,435),(34,430),(35,431),(36,432),(37,423),(38,422),(39,421),(40,428),(41,429),
(42,425),(43,424),(44,427),(45,426),(46,484),(47,485),(48,486),(49,487),(50,488),(51,489),
(52,490),(53,491),(54,492),(55,493),(56,494),(57,495),(58,445),(59,446),(60,447),(61,438),
(62,437),(63,436),(64,443),(65,444),(66,440),(67,439),(68,442),(69,441),(70,457),(71,458),
(72,459),(73,450),(74,449),(75,448),(76,455),(77,456),(78,452),(79,451),(80,454),(81,453),
(82,478),(83,479),(84,480),(85,481),(86,482),(87,483),(88,464),(89,465),(90,461),(91,460),
(92,463),(93,462),(94,470),(95,471),(96,467),(97,466),(98,469),(99,468),(100,476),
(101,477),(102,473),(103,472),(104,475),(105,474),(106,1021),(107,1022),(108,1023),
(109,1024),(110,1025),(111,1026),(112,1027),(113,1028),(114,1029),(115,1030),(116,1031),
(117,1032),(118,1033),(119,1034),(120,1035),(121,1036),(122,1037),(123,1038),(124,1039),
(125,1040),(126,1041),(127,1042),(128,1043),(129,1044),(130,1045),(131,1046),(132,1047),
(133,1048),(134,1049),(135,1050),(136,1051),(137,1052),(138,1053),(139,1054),(140,1055),
(141,1056),(142,1057),(143,1058),(144,1059),(145,1060),(146,1061),(147,1062),(148,1063),
(149,1064),(150,1065),(151,931),(152,932),(153,933),(154,934),(155,935),(156,936),
(157,937),(158,938),(159,939),(160,940),(161,941),(162,942),(163,943),(164,944),(165,945),
(166,946),(167,947),(168,948),(169,949),(170,950),(171,951),(172,952),(173,953),(174,954),
(175,955),(176,956),(177,957),(178,958),(179,959),(180,960),(181,853),(182,854),(183,855),
(184,856),(185,857),(186,858),(187,844),(188,845),(189,846),(190,841),(191,842),(192,843),
(193,849),(194,850),(195,847),(196,848),(197,852),(198,851),(199,865),(200,866),(201,867),
(202,861),(203,862),(204,859),(205,860),(206,864),(207,863),(208,870),(209,869),(210,868),
(211,895),(212,896),(213,897),(214,898),(215,899),(216,900),(217,901),(218,902),(219,903),
(220,904),(221,905),(222,906),(223,874),(224,875),(225,876),(226,871),(227,872),(228,873),
(229,879),(230,880),(231,877),(232,878),(233,882),(234,881),(235,886),(236,887),(237,888),
(238,883),(239,884),(240,885),(241,891),(242,892),(243,889),(244,890),(245,894),(246,893),
(247,919),(248,920),(249,921),(250,922),(251,923),(252,924),(253,909),(254,910),(255,907),
(256,908),(257,912),(258,911),(259,915),(260,916),(261,913),(262,914),(263,918),(264,917),
(265,929),(266,930),(267,926),(268,925),(269,928),(270,927),(271,997),(272,998),(273,999),
(274,1000),(275,1001),(276,1002),(277,1003),(278,1004),(279,1005),(280,1006),(281,1007),
(282,1008),(283,967),(284,968),(285,969),(286,963),(287,964),(288,961),(289,962),
(290,966),(291,965),(292,972),(293,971),(294,970),(295,985),(296,986),(297,987),(298,988),
(299,989),(300,990),(301,975),(302,976),(303,973),(304,974),(305,978),(306,977),(307,981),
(308,982),(309,979),(310,980),(311,984),(312,983),(313,995),(314,996),(315,992),(316,991),
(317,994),(318,993),(319,1018),(320,1019),(321,1020),(322,1011),(323,1010),(324,1009),
(325,1016),(326,1017),(327,1013),(328,1012),(329,1015),(330,1014),(331,1126),(332,1127),
(333,1128),(334,1129),(335,1130),(336,1131),(337,1132),(338,1133),(339,1134),(340,1135),
(341,1136),(342,1137),(343,1138),(344,1139),(345,1140),(346,1102),(347,1103),(348,1104),
(349,1105),(350,1106),(351,1107),(352,1108),(353,1109),(354,1110),(355,$1$),(356,1112),
(357,1113),(358,1072),(359,1073),(360,1074),(361,1068),(362,1069),(363,1066),(364,1067),
(365,1071),(366,1070),(367,1077),(368,1076),(369,1075),(370,1090),(371,1091),(372,1092),
(373,1093),(374,1094),(375,1095),(376,1080),(377,1081),(378,1078),(379,1079),(380,1083),
(381,1082),(382,1086),(383,1087),(384,1084),(385,1085),(386,1089),(387,1088),(388,1100),
(389,1101),(390,1097),(391,1096),(392,1099),(393,1098),(394,1123),(395,1124),(396,1125),
(397,1116),(398,1115),(399,1114),(400,1121),(401,1122),(402,1118),(403,1117),(404,1120),
(405,1119),(406,1153),(407,1154),(408,1155),(409,1150),(410,1151),(411,1152),(412,1143),
(413,1142),(414,1141),(415,1148),(416,1149),(417,1145),(418,1144),(419,1147),(420,1146)
Colour 1:
(1,511),(2,512),(3,513),(4,514),(5,515),(6,516),(7,517),(8,518),(9,519),(10,520),(11,521),
(12,522),(13,523),(14,524),(15,525),(16,31),(17,32),(18,33),(19,34),(20,35),(21,36),
(22,37),(23,38),(24,39),(25,40),(26,41),(27,42),(28,43),(29,44),(30,45),(46,58),(47,59),
(48,60),(49,61),(50,62),(51,63),(52,64),(53,65),(54,66),(55,67),(56,68),(57,69),(70,73),
(71,74),(72,75),(76,78),(77,79),(80,81),(82,88),(83,89),(84,90),(85,91),(86,92),(87,93),
(94,96),(95,97),(98,99),(100,102),(101,103),(104,105),(106,1021),(107,1022),(108,1023),
(109,1024),(110,1025),(111,1026),(112,1027),(113,1028),(114,1029),(115,1030),(116,1031),
(117,1032),(118,1033),(119,1034),(120,1035),(121,1036),(122,1037),(123,1038),(124,1039),
(125,1040),(126,1041),(127,1042),(128,1043),(129,1044),(130,1045),(131,1046),(132,1047),
(133,1048),(134,1049),(135,1050),(136,1051),(137,1052),(138,1053),(139,1054),(140,1055),
(141,1056),(142,1057),(143,1058),(144,1059),(145,1060),(146,1061),(147,1062),(148,1063),
(149,1064),(150,1065),(151,181),(152,182),(153,183),(154,184),(155,185),(156,186),
(157,187),(158,188),(159,189),(160,190),(161,191),(162,192),(163,193),(164,194),(165,195),
(166,196),(167,197),(168,198),(169,199),(170,200),(171,201),(172,202),(173,203),(174,204),
(175,205),(176,206),(177,207),(178,208),(179,209),(180,210),(211,223),(212,224),(213,225),
(214,226),(215,227),(216,228),(217,229),(218,230),(219,231),(220,232),(221,233),(222,234),
(235,238),(236,239),(237,240),(241,243),(242,244),(245,246),(247,253),(248,254),(249,255),
(250,256),(251,257),(252,258),(259,261),(260,262),(263,264),(265,267),(266,268),(269,270),
(271,283),(272,284),(273,285),(274,286),(275,287),(276,288),(277,289),(278,290),(279,291),
(280,292),(281,293),(282,294),(295,301),(296,302),(297,303),(298,304),(299,305),(300,306),
(307,309),(308,310),(311,312),(313,315),(314,316),(317,318),(319,322),(320,323),(321,324),
(325,327),(326,328),(329,330),(331,421),(332,422),(333,423),(334,424),(335,425),(336,426),
(337,427),(338,428),(339,429),(340,430),(341,431),(342,432),(343,433),(344,434),(345,435),
(346,436),(347,437),(348,438),(349,439),(350,440),(351,441),(352,442),(353,443),(354,444),
(355,445),(356,446),(357,447),(358,448),(359,449),(360,450),(361,451),(362,452),(363,453),
(364,454),(365,455),(366,456),(367,457),(368,458),(369,459),(370,460),(371,461),(372,462),
(373,463),(374,464),(375,465),(376,466),(377,467),(378,468),(379,469),(380,470),(381,471),
(382,472),(383,473),(384,474),(385,475),(386,476),(387,477),(388,478),(389,479),(390,480),
(391,481),(392,482),(393,483),(394,484),(395,485),(396,486),(397,487),(398,488),(399,489),
(400,490),(401,491),(402,492),(403,493),(404,494),(405,495),(406,496),(407,497),(408,498),
(409,499),(410,500),(411,501),(412,502),(413,503),(414,504),(415,505),(416,506),(417,507),
(418,508),(419,509),(420,510),(571,586),(572,587),(573,588),(574,589),(575,590),(576,591),
(577,592),(578,593),(579,594),(580,595),(581,596),(582,597),(583,598),(584,599),(585,600),
(601,613),(602,614),(603,615),(604,616),(605,617),(606,618),(607,619),(608,620),(609,621),
(610,622),(611,623),(612,624),(625,628),(626,629),(627,630),(631,633),(632,634),(635,636),
(637,643),(638,644),(639,645),(640,646),(641,647),(642,648),(649,651),(650,652),(653,654),
(655,657),(656,658),(659,660),(661,841),(662,842),(663,843),(664,844),(665,845),(666,846),
(667,847),(668,848),(669,849),(670,850),(671,851),(672,852),(673,853),(674,854),(675,855),
(676,856),(677,857),(678,858),(679,859),(680,860),(681,861),(682,862),(683,863),(684,864),
(685,865),(686,866),(687,867),(688,868),(689,869),(690,870),(691,871),(692,872),(693,873),
(694,874),(695,875),(696,876),(697,877),(698,878),(699,879),(700,880),(701,881),(702,882),
(703,883),(704,884),(705,885),(706,886),(707,887),(708,888),(709,889),(710,890),(711,891),
(712,892),(713,893),(714,894),(715,895),(716,896),(717,897),(718,898),(719,899),(720,900),
(721,901),(722,902),(723,903),(724,904),(725,905),(726,906),(727,907),(728,908),(729,909),
(730,910),(731,911),(732,912),(733,913),(734,914),(735,915),(736,916),(737,917),(738,918),
(739,919),(740,920),(741,921),(742,922),(743,923),(744,924),(745,925),(746,926),(747,927),
(748,928),(749,929),(750,930)
Colour 2:
(1,16),(2,17),(3,18),(4,19),(5,20),(6,21),(7,22),(8,23),(9,24),(10,25),(11,26),(12,27),
(13,28),(14,29),(15,30),(31,406),(32,407),(33,408),(34,409),(35,410),(36,411),(37,412),
(38,413),(39,414),(40,415),(41,416),(42,417),(43,418),(44,419),(45,420),(46,49),(47,50),
(48,51),(52,54),(53,55),(56,57),(58,70),(59,71),(60,72),(61,73),(62,74),(63,75),(64,76),
(65,77),(66,78),(67,79),(68,80),(69,81),(82,84),(83,85),(86,87),(88,94),(89,95),(90,96),
(91,97),(92,98),(93,99),(100,101),(102,104),(103,105),(106,151),(107,152),(108,153),
(109,154),(110,155),(111,156),(112,157),(113,158),(114,159),(115,160),(116,161),(117,162),
(118,163),(119,164),(120,165),(121,166),(122,167),(123,168),(124,169),(125,170),(126,171),
(127,172),(128,173),(129,174),(130,175),(131,176),(132,177),(133,178),(134,179),(135,180),
(136,331),(137,332),(138,333),(139,334),(140,335),(141,336),(142,337),(143,338),(144,339),
(145,340),(146,341),(147,342),(148,343),(149,344),(150,345),(181,751),(182,752),(183,753),
(184,754),(185,755),(186,756),(187,757),(188,758),(189,759),(190,760),(191,761),(192,762),
(193,763),(194,764),(195,765),(196,766),(197,767),(198,768),(199,769),(200,770),(201,771),
(202,772),(203,773),(204,774),(205,775),(206,776),(207,777),(208,778),(209,779),(210,780),
(211,214),(212,215),(213,216),(217,219),(218,220),(221,222),(223,235),(224,236),(225,237),
(226,238),(227,239),(228,240),(229,241),(230,242),(231,243),(232,244),(233,245),(234,246),
(247,249),(248,250),(251,252),(253,259),(254,260),(255,261),(256,262),(257,263),(258,264),
(265,266),(267,269),(268,270),(271,346),(272,347),(273,348),(274,349),(275,350),(276,351),
(277,352),(278,353),(279,354),(280,355),(281,356),(282,357),(283,358),(284,359),(285,360),
(286,361),(287,362),(288,363),(289,364),(290,365),(291,366),(292,367),(293,368),(294,369),
(295,370),(296,371),(297,372),(298,373),(299,374),(300,375),(301,376),(302,377),(303,378),
(304,379),(305,380),(306,381),(307,382),(308,383),(309,384),(310,385),(311,386),(312,387),
(313,388),(314,389),(315,390),(316,391),(317,392),(318,393),(319,394),(320,395),(321,396),
(322,397),(323,398),(324,399),(325,400),(326,401),(327,402),(328,403),(329,404),(330,405),
(421,1141),(422,1142),(423,1143),(424,1144),(425,1145),(426,1146),(427,1147),(428,1148),
(429,1149),(430,1150),(431,1151),(432,1152),(433,1153),(434,1154),(435,1155),(436,448),
(437,449),(438,450),(439,451),(440,452),(441,453),(442,454),(443,455),(444,456),(445,457),
(446,458),(447,459),(460,466),(461,467),(462,468),(463,469),(464,470),(465,471),(472,474),
(473,475),(476,477),(478,480),(479,481),(482,483),(484,487),(485,488),(486,489),(490,492),
(491,493),(494,495),(496,511),(497,512),(498,513),(499,514),(500,515),(501,516),(502,517),
(503,518),(504,519),(505,520),(506,521),(507,522),(508,523),(509,524),(510,525),(526,571),
(527,572),(528,573),(529,574),(530,575),(531,576),(532,577),(533,578),(534,579),(535,580),
(536,581),(537,582),(538,583),(539,584),(540,585),(541,661),(542,662),(543,663),(544,664),
(545,665),(546,666),(547,667),(548,668),(549,669),(550,670),(551,671),(552,672),(553,673),
(554,674),(555,675),(556,676),(557,677),(558,678),(559,679),(560,680),(561,681),(562,682),
(563,683),(564,684),(565,685),(566,686),(567,687),(568,688),(569,689),(570,690),(601,691),
(602,692),(603,693),(604,694),(605,695),(606,696),(607,697),(608,698),(609,699),(610,700),
(611,701),(612,702),(613,703),(614,704),(615,705),(616,706),(617,707),(618,708),(619,709),
(620,710),(621,711),(622,712),(623,713),(624,714),(625,715),(626,716),(627,717),(628,718),
(629,719),(630,720),(631,721),(632,722),(633,723),(634,724),(635,725),(636,726),(637,727),
(638,728),(639,729),(640,730),(641,731),(642,732),(643,733),(644,734),(645,735),(646,736),
(647,737),(648,738),(649,739),(650,740),(651,741),(652,742),(653,743),(654,744),(655,745),
(656,746),(657,747),(658,748),(659,749),(660,750),(781,782),(783,785),(784,786),(787,790),
(788,791),(789,792),(793,794),(795,797),(796,798),(799,800),(801,803),(802,804),(805,811),
(806,812),(807,813),(808,814),(809,815),(810,816),(817,829),(818,830),(819,831),(820,832),
(821,833),(822,834),(823,835),(824,836),(825,837),(826,838),(827,839),(828,840)
Colour 3:
(1,148),(2,149),(3,150),(4,7),(5,8),(6,9),(10,12),(11,13),(14,15),(16,31),(17,32),(18,33),
(19,46),(20,47),(21,48),(22,49),(23,50),(24,51),(25,52),(26,53),(27,54),(28,55),(29,56),
(30,57),(34,58),(35,59),(36,60),(37,61),(38,62),(39,63),(40,64),(41,65),(42,66),(43,67),
(44,68),(45,69),(70,319),(71,320),(72,321),(73,322),(74,323),(75,324),(76,325),(77,326),
(78,327),(79,328),(80,329),(81,330),(82,83),(84,86),(85,87),(88,89),(90,92),(91,93),
(94,100),(95,101),(96,102),(97,103),(98,104),(99,105),(106,553),(107,554),(108,555),
(109,556),(110,557),(111,558),(112,115),(113,116),(114,117),(118,120),(119,121),(122,123),
(124,136),(125,137),(126,138),(127,139),(128,140),(129,141),(130,142),(131,143),(132,144),
(133,145),(134,146),(135,147),(151,181),(152,182),(153,183),(154,184),(155,185),(156,186),
(157,211),(158,212),(159,213),(160,214),(161,215),(162,216),(163,217),(164,218),(165,219),
(166,220),(167,221),(168,222),(169,271),(170,272),(171,273),(172,274),(173,275),(174,276),
(175,277),(176,278),(177,279),(178,280),(179,281),(180,282),(187,223),(188,224),(189,225),
(190,226),(191,227),(192,228),(193,229),(194,230),(195,231),(196,232),(197,233),(198,234),
(199,283),(200,284),(201,285),(202,286),(203,287),(204,288),(205,289),(206,290),(207,291),
(208,292),(209,293),(210,294),(235,625),(236,626),(237,627),(238,628),(239,629),(240,630),
(241,631),(242,632),(243,633),(244,634),(245,635),(246,636),(247,295),(248,296),(249,297),
(250,298),(251,299),(252,300),(253,301),(254,302),(255,303),(256,304),(257,305),(258,306),
(259,307),(260,308),(261,309),(262,310),(263,311),(264,312),(265,313),(266,314),(267,315),
(268,316),(269,317),(270,318),(331,346),(332,347),(333,348),(334,349),(335,350),(336,351),
(337,352),(338,353),(339,354),(340,355),(341,356),(342,357),(343,406),(344,407),(345,408),
(358,829),(359,830),(360,831),(361,832),(362,833),(363,834),(364,835),(365,836),(366,837),
(367,838),(368,839),(369,840),(370,372),(371,373),(374,375),(376,382),(377,383),(378,384),
(379,385),(380,386),(381,387),(388,389),(390,392),(391,393),(394,409),(395,410),(396,411),
(397,412),(398,413),(399,414),(400,415),(401,416),(402,417),(403,418),(404,419),(405,420),
(421,436),(422,437),(423,438),(424,439),(425,440),(426,441),(427,442),(428,443),(429,444),
(430,445),(431,446),(432,447),(433,496),(434,497),(435,498),(448,1009),(449,1010),
(450,1011),(451,1012),(452,1013),(453,1014),(454,1015),(455,1016),(456,1017),(457,1018),
(458,1019),(459,1020),(460,462),(461,463),(464,465),(466,472),(467,473),(468,474),
(469,475),(470,476),(471,477),(478,479),(480,482),(481,483),(484,499),(485,500),(486,501),
(487,502),(488,503),(489,504),(490,505),(491,506),(492,507),(493,508),(494,509),(495,510),
(511,1063),(512,1064),(513,1065),(514,517),(515,518),(516,519),(520,522),(521,523),
(524,525),(529,541),(530,542),(531,543),(532,544),(533,545),(534,546),(535,547),(536,548),
(537,549),(538,550),(539,551),(540,552),(559,560),(561,563),(562,564),(565,568),(566,569),
(567,570),(571,586),(572,587),(573,588),(574,601),(575,602),(576,603),(577,604),(578,605),
(579,606),(580,607),(581,608),(582,609),(583,610),(584,611),(585,612),(589,613),(590,614),
(591,615),(592,616),(593,617),(594,618),(595,619),(596,620),(597,621),(598,622),(599,623),
(600,624),(637,638),(639,641),(640,642),(643,644),(645,647),(646,648),(649,655),(650,656),
(651,657),(652,658),(653,659),(654,660),(661,691),(662,692),(663,693),(664,694),(665,695),
(666,696),(667,697),(668,698),(669,699),(670,700),(671,701),(672,702),(673,751),(674,752),
(675,753),(676,754),(677,755),(678,756),(679,781),(680,782),(681,783),(682,784),(683,785),
(684,786),(685,787),(686,788),(687,789),(688,790),(689,791),(690,792),(715,757),(716,758),
(717,759),(718,760),(719,761),(720,762),(721,763),(722,764),(723,765),(724,766),(725,767),
(726,768),(727,793),(728,794),(729,795),(730,796),(731,797),(732,798),(733,799),(734,800),
(735,801),(736,802),(737,803),(738,804),(739,805),(740,806),(741,807),(742,808),(743,809),
(744,810),(745,811),(746,812),(747,813),(748,814),(749,815),(750,816),(769,817),(770,818),
(771,819),(772,820),(773,821),(774,822),(775,823),(776,824),(777,825),(778,826),(779,827),
(780,828)
Colour 4:
(1,4),(2,5),(3,6),(7,133),(8,134),(9,135),(10,11),(12,14),(13,15),(16,19),(17,20),(18,21),
(22,178),(23,179),(24,180),(25,26),(27,29),(28,30),(31,34),(32,35),(33,36),(37,208),
(38,209),(39,210),(40,41),(42,44),(43,45),(46,58),(47,59),(48,60),(49,70),(50,71),(51,72),
(52,82),(53,83),(54,84),(55,85),(56,86),(57,87),(61,73),(62,74),(63,75),(64,88),(65,89),
(66,90),(67,91),(68,92),(69,93),(76,94),(77,95),(78,96),(79,97),(80,98),(81,99),(100,265),
(101,266),(102,267),(103,268),(104,269),(105,270),(106,112),(107,113),(108,114),(109,124),
(110,125),(111,126),(115,532),(116,533),(117,534),(118,127),(119,128),(120,129),(121,130),
(122,131),(123,132),(136,568),(137,569),(138,570),(139,141),(140,142),(143,144),(145,148),
(146,149),(147,150),(151,157),(152,158),(153,159),(154,169),(155,170),(156,171),(160,577),
(161,578),(162,579),(163,172),(164,173),(165,174),(166,175),(167,176),(168,177),(181,187),
(182,188),(183,189),(184,199),(185,200),(186,201),(190,592),(191,593),(192,594),(193,202),
(194,203),(195,204),(196,205),(197,206),(198,207),(211,223),(212,224),(213,225),(214,235),
(215,236),(216,237),(217,247),(218,248),(219,249),(220,250),(221,251),(222,252),(226,238),
(227,239),(228,240),(229,253),(230,254),(231,255),(232,256),(233,257),(234,258),(241,259),
(242,260),(243,261),(244,262),(245,263),(246,264),(271,283),(272,284),(273,285),(274,295),
(275,296),(276,297),(277,298),(278,299),(279,300),(280,319),(281,320),(282,321),(286,301),
(287,302),(288,303),(289,304),(290,305),(291,306),(292,322),(293,323),(294,324),(307,655),
(308,656),(309,657),(310,658),(311,659),(312,660),(313,325),(314,326),(315,327),(316,328),
(317,329),(318,330),(331,688),(332,689),(333,690),(334,336),(335,337),(338,339),(340,343),
(341,344),(342,345),(346,358),(347,359),(348,360),(349,370),(350,371),(351,372),(352,373),
(353,374),(354,375),(355,394),(356,395),(357,396),(361,376),(362,377),(363,378),(364,379),
(365,380),(366,381),(367,397),(368,398),(369,399),(382,745),(383,746),(384,747),(385,748),
(386,749),(387,750),(388,400),(389,401),(390,402),(391,403),(392,404),(393,405),(406,409),
(407,410),(408,411),(412,778),(413,779),(414,780),(415,416),(417,419),(418,420),(421,868),
(422,869),(423,870),(424,426),(425,427),(428,429),(430,433),(431,434),(432,435),(436,448),
(437,449),(438,450),(439,460),(440,461),(441,462),(442,463),(443,464),(444,465),(445,484),
(446,485),(447,486),(451,466),(452,467),(453,468),(454,469),(455,470),(456,471),(457,487),
(458,488),(459,489),(472,925),(473,926),(474,927),(475,928),(476,929),(477,930),(478,490),
(479,491),(480,492),(481,493),(482,494),(483,495),(496,499),(497,500),(498,501),(502,958),
(503,959),(504,960),(505,506),(507,509),(508,510),(511,514),(512,515),(513,516),
(517,1048),(518,1049),(519,1050),(520,521),(522,524),(523,525),(526,529),(527,530),
(528,531),(535,536),(537,539),(538,540),(544,553),(545,554),(546,555),(547,559),(548,560),
(549,561),(550,562),(551,563),(552,564),(556,565),(557,566),(558,567),(571,574),(572,575),
(573,576),(580,581),(582,584),(583,585),(586,589),(587,590),(588,591),(595,596),(597,599),
(598,600),(601,613),(602,614),(603,615),(604,625),(605,626),(606,627),(607,637),(608,638),
(609,639),(610,640),(611,641),(612,642),(616,628),(617,629),(618,630),(619,643),(620,644),
(621,645),(622,646),(623,647),(624,648),(631,649),(632,650),(633,651),(634,652),(635,653),
(636,654),(664,673),(665,674),(666,675),(667,679),(668,680),(669,681),(670,682),(671,683),
(672,684),(676,685),(677,686),(678,687),(691,703),(692,704),(693,705),(694,715),(695,716),
(696,717),(697,727),(698,728),(699,729),(700,730),(701,731),(702,732),(706,718),(707,719),
(708,720),(709,733),(710,734),(711,735),(712,736),(713,737),(714,738),(721,739),(722,740),
(723,741),(724,742),(725,743),(726,744),(751,757),(752,758),(753,759),(754,769),(755,770),
(756,771),(763,772),(764,773),(765,774),(766,775),(767,776),(768,777),(781,793),(782,794),
(783,795),(784,796),(785,797),(786,798),(787,817),(788,818),(789,819),(790,829),(791,830),
(792,831),(805,820),(806,821),(807,822),(808,823),(809,824),(810,825),(811,832),(812,833),
(813,834),(814,835),(815,836),(816,837),(826,838),(827,839),(828,840)